GNU Generic Security Service Library

GSS-API Library for the GNU system for version 1.0.3, 9 October 2014

Simon Josefsson

This manual is last updated 9 October 2014 for version 1.0.3 of GNU GSS. Copyright \begin{matrix}c \\ ◯ \end{matrix} 2003-2014 Simon Josefsson.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Getting Started . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 GSS-API Overview. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.4 Supported Platforms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.5 Commercial Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.6 Downloading and Installing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.7 Bug Reports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.8 Contributing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.9 Planned Features. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

2 Preparation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.1 Header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2 Initialization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.3 Version Check. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.4 Building the source. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.5 Out of Memory handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

3 Standard GSS API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

3.1 Simple Data Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.1.1 Integer types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.1.2 String and similar data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

3.1.2.1 Opaque data types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.1.2.2 Character strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

3.1.3 Object Identifiers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.1.4 Object Identifier Sets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

3.2 Complex Data Types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.2.1 Credentials. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.2.2 Contexts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.2.3 Authentication tokens . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.2.4 Interprocess tokens . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.2.5 Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.2.6 Channel Bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

3.3 Optional Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.4 Error Handling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

3.4.1 GSS status codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.4.2 Mechanism-specific status codes . . . . . . . . . . . . . . . . . . . . . . . . . . 18

3.5 Credential Management . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.6 Context-Level Routines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.7 Per-Message Routines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 3.8 Name Manipulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 3.9 Miscellaneous Routines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 3.10 SASL GS2 Routines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

4 Extended GSS API. . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5 Invoking gss. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 6 Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 Appendix A Criticism of GSS. . . . . . . . . . . . . . . . . . 61 Appendix B Copying Information . . . . . . . . . . . . . 63

B.1 GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 API Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

1 Introduction

GSS is an implementation of the Generic Security Service Application Program Interface (GSS-API). GSS-API is used by network servers to provide security services, e.g., to authenticate SMTP/IMAP clients against SMTP/IMAP servers. GSS consists of a library and a manual.

GSS is developed for the GNU/Linux system, but runs on over 20 platforms including most major Unix platforms and Windows, and many kind of devices including iPAQ handhelds and S/390 mainframes.

GSS is a GNU project, and is licensed under the GNU General Public License version 3 or later.

1.1 Getting Started

This manual documents the GSS programming interface. All functions and data types provided by the library are explained.

The reader is assumed to possess basic familiarity with GSS-API and network programming in C or C++. For general GSS-API information, and some programming examples,

there is a guide available online at http://docs.sun.com/db/doc/816-1331.

This manual can be used in several ways. If read from the beginning to the end, it gives a good introduction into the library and how it can be used in an application. Forward references are included where necessary. Later on, the manual can be used as a reference manual to get just the information needed about any particular interface of the library. Experienced programmers might want to start looking at the examples at the end of the manual, and then only read up those parts of the interface which are unclear.

1.2 Features

GSS might have a couple of advantages over other libraries doing a similar job.

It’s Free Software Anybody can use, modify, and redistribute it under the terms of the GNU General Public License version 3 or later.

It’s thread-safe No global variables are used and multiple library handles and session handles may be used in parallell.

It’s internationalized It handles non-ASCII names and user visible strings used in the library (e.g., error messages) can be translated into the users’ language.

It’s portable It should work on all Unix like operating systems, including Windows.

1.3 GSS-API Overview

This section describes GSS-API from a protocol point of view.

The Generic Security Service Application Programming Interface provides security services to calling applications. It allows a communicating application to authenticate the user associated with another application, to delegate rights to another application, and to apply security services such as confidentiality and integrity on a per-message basis.

There are four stages to using the GSS-API:

1. The application acquires a set of credentials with which it may prove its identity to other processes. The application’s credentials vouch for its global identity, which may or may not be related to any local username under which it may be running.

2. A pair of communicating applications establish a joint security context using their credentials. The security context is a pair of GSS-API data structures that contain shared state information, which is required in order that per-message security services may be provided. Examples of state that might be shared between applications as part of a security context are cryptographic keys, and message sequence numbers. As part of the establishment of a security context, the context initiator is authenticated to the responder, and may require that the responder is authenticated in turn. The initiator may optionally give the responder the right to initiate further security contexts, acting as an agent or delegate of the initiator. This transfer of rights is termed delegation, and is achieved by creating a set of credentials, similar to those used by the initiating application, but which may be used by the responder. To establish and maintain the shared information that makes up the security context, certain GSS-API calls will return a token data structure, which is an opaque data type that may contain cryptographically protected data. The caller of such a GSS-API routine is responsible for transferring the token to the peer application, encapsulated if necessary in an application- application protocol. On receipt of such a token, the peer application should pass it to a corresponding GSS-API routine which will decode the token and extract the information, updating the security context state information accordingly.

3. Per-message services are invoked to apply either: integrity and data origin authentication, or confidentiality, integrity and data origin authentication to application data, which are treated by GSS-API as arbitrary octet-strings. An application transmitting a message that it wishes to protect will call the appropriate GSS-API routine (gss get mic or gss wrap) to apply protection, specifying the appropriate security context, and send the resulting token to the receiving application. The receiver will pass the received token (and, in the case of data protected by gss get mic, the accompanying message-data) to the corresponding decoding routine (gss verify mic or gss unwrap) to remove the protection and validate the data.

4. At the completion of a communications session (which may extend across several transport connections), each application calls a GSS-API routine to delete the security context. Multiple contexts may also be used (either successively or simultaneously) within a single communications association, at the option of the applications.

1.4 Supported Platforms

GSS has at some point in time been tested on the following platforms.

1. Debian GNU/Linux 3.0 (Woody)

GCC 2.95.4 and GNU Make. This is the main development platform.

alphaev67-unknown-linux-gnu, alphaev6-unknown-linux-gnu, arm-unknown-
linux-gnu, hppa-unknown-linux-gnu, hppa64-unknown-linux-gnu, i686-pc-

linux-gnu, ia64-unknown-linux-gnu, m68k-unknown-linux-gnu, mips-unknown-
linux-gnu,  mipsel-unknown-linux-gnu,  powerpc-unknown-linux-gnu,
s390-ibm-linux-gnu, sparc-unknown-linux-gnu.

2. Debian GNU/Linux 2.1

GCC 2.95.1 and GNU Make. armv4l-unknown-linux-gnu.

3. Tru64 UNIX

Tru64 UNIX C compiler and Tru64 Make. alphaev67-dec-osf5.1, alphaev68-dec-
osf5.1.

4. SuSE Linux 7.1

GCC 2.96 and GNU Make. alphaev6-unknown-linux-gnu, alphaev67-unknown-
linux-gnu.

5. SuSE Linux 7.2a

GCC 3.0 and GNU Make. ia64-unknown-linux-gnu.

6. RedHat Linux 7.2

GCC 2.96 and GNU Make. alphaev6-unknown-linux-gnu, alphaev67-unknown-
linux-gnu, ia64-unknown-linux-gnu.

7. RedHat Linux 8.0 GCC 3.2 and GNU Make. i686-pc-linux-gnu.

8. RedHat Advanced Server 2.1 GCC 2.96 and GNU Make. i686-pc-linux-gnu.

9. Slackware Linux 8.0.01 GCC 2.95.3 and GNU Make. i686-pc-linux-gnu.

10. Mandrake Linux 9.0 GCC 3.2 and GNU Make. i686-pc-linux-gnu.

11. IRIX 6.5 MIPS C compiler, IRIX Make. mips-sgi-irix6.5.

12. AIX 4.3.2 IBM C for AIX compiler, AIX Make. rs6000-ibm-aix4.3.2.0.

13. Microsoft Windows 2000 (Cygwin) GCC 3.2, GNU make. i686-pc-cygwin.

14. HP-UX 11

HP-UX C compiler and HP Make. ia64-hp-hpux11.22, hppa2.0w-hp-hpux11.11.

15. SUN Solaris 2.8 Sun WorkShop Compiler C 6.0 and SUN Make. sparc-sun-solaris2.8.

16. NetBSD 1.6

GCC 2.95.3 and GNU Make.  alpha-unknown-netbsd1.6, i386-unknown-
netbsdelf1.6.

17. OpenBSD 3.1 and 3.2

GCC 2.95.3 and GNU Make.  alpha-unknown-openbsd3.1, i386-unknown-
openbsd3.1.

18. FreeBSD 4.7

    GCC 2.95.4 and GNU Make.  alpha-unknown-freebsd4.7, i386-unknown-
    freebsd4.7.

19. Cross compiled to uClinux/uClibc on Motorola Coldfire. GCC 3.4 and GNU Make m68k-uclinux-elf.

If you use GSS on, or port GSS to, a new platform please report it to the author.

1.5 Commercial Support

Commercial support is available for users of GNU GSS. The kind of support that can be purchased may include:

• Implement new features. Such as a new GSS-API mechanism.

• Port GSS to new platforms. This could include porting to an embedded platforms that may need memory or size optimization.

• Integrating GSS as a security environment in your existing project.

• System design of components related to GSS-API.

If you are interested, please write to:

Simon Josefsson Datakonsult AB
Hagagatan 24
113 47 Stockholm
Sweden
E-mail: simon@josefsson.org

If your company provides support related to GNU GSS and would like to be mentioned here, contact the author (see Section 1.7 [Bug Reports], page 5).

1.6 Downloading and Installing

The package can be downloaded from several places, including:

   ftp://ftp.gnu.org/gnu/gss/

The latest version is stored in a file, e.g., ‘gss-1.0.3.tar.gz’ where the ‘1.0.3’ indicate the highest version number.

The package is then extracted, configured and built like many other packages that use Autoconf. For detailed information on configuring and building it, refer to the INSTALL file that is part of the distribution archive.

Here is an example terminal session that downloads, configures, builds and installs the package. You will need a few basic tools, such as ‘sh’, ‘make’ and ‘cc’.

     $ wget -q ftp://ftp.gnu.org/gnu/gss/gss-1.0.3.tar.gz
     $ tar xfz gss-1.0.3.tar.gz
     $ cd gss-1.0.3/
     $ ./configure
     ...
     $ make
     ...

     $ make install
     ...

After that GSS should be properly installed and ready for use.

1.7 Bug Reports

If you think you have found a bug in GSS, please investigate it and report it.

• Please make sure that the bug is really in GSS, and preferably also check that it hasn’t already been fixed in the latest version.

• You have to send us a test case that makes it possible for us to reproduce the bug.

• You also have to explain what is wrong; if you get a crash, or if the results printed are not good and in that case, in what way. Make sure that the bug report includes all information you would need to fix this kind of bug for someone else.

Please make an effort to produce a self-contained report, with something definite that can be tested or debugged. Vague queries or piecemeal messages are difficult to act on and don’t help the development effort.

If your bug report is good, we will do our best to help you to get a corrected version of the software; if the bug report is poor, we won’t do anything about it (apart from asking you to send better bug reports).

If you think something in this manual is unclear, or downright incorrect, or if the language needs to be improved, please also send a note.

Send your bug report to:

                              ‘bug-gss@gnu.org’

1.8 Contributing

If you want to submit a patch for inclusion – from solve a typo you discovered, up to adding support for a new feature – you should submit it as a bug report (see Section 1.7 [Bug Reports], page 5). There are some things that you can do to increase the chances for it to be included in the official package.

Unless your patch is very small (say, under 10 lines) we require that you assign the copyright of your work to the Free Software Foundation. This is to protect the freedom of the project. If you have not already signed papers, we will send you the necessary information when you submit your contribution.

For contributions that doesn’t consist of actual programming code, the only guidelines are common sense. Use it.

For code contributions, a number of style guides will help you:

• Coding Style. Follow the GNU Standards document (see Section “top” in standards). If you normally code using another coding standard, there is no problem, but you should use ‘indent’ to reformat the code (see Section “top” in indent) before submitting your work.

• Use the unified diff format ‘diff -u’.

• Return errors. No reason whatsoever should abort the execution of the library. Even memory allocation errors, e.g. when malloc return NULL, should work although result in an error code.

• Design with thread safety in mind. Don’t use global variables. Don’t even write to per-handle global variables unless the documented behaviour of the function you write is to write to the per-handle global variable.

• Avoid using the C math library. It causes problems for embedded implementations, and in most situations it is very easy to avoid using it.

• Document your functions. Use comments before each function headers, that, if properly formatted, are extracted into Texinfo manuals and GTK-DOC web pages.

• Supply a ChangeLog and NEWS entries, where appropriate.

1.9 Planned Features

This is also known as the “todo list”. If you like to start working on anything, please let me know so work duplication can be avoided.

• Support non-blocking mode. This would be an API extension. It could work by forking a process and interface to it, or by using a user-specific daemon. E.g., h = START(accept sec context(...)), FINISHED(h), ret = FINISH(h), ABORT(h).

• Support loadable modules via dlopen, a’la Solaris GSS.

• Port to Cyclone? CCured?

2 Preparation

To use GSS, you have to perform some changes to your sources and the build system. The necessary changes are small and explained in the following sections. At the end of this chapter, it is described how the library is initialized, and how the requirements of the library are verified.

A faster way to find out how to adapt your application for use with GSS may be to look at the examples at the end of this manual.

2.1 Header

All standard interfaces (data types and functions) of the official GSS API are defined in the header file gss/api.h. The file is taken verbatim from the RFC (after correcting a few typos) where it is known as gssapi.h. However, to be able to co-exist gracefully with other GSS-API implementation, the name gssapi.h was changed.

The header file gss.h includes gss/api.h, and declares a few non-standard extensions (by including gss/ext.h), takes care of including header files related to all supported mechanisms (e.g., gss/krb5.h) and finally adds C++ namespace protection of all definitions. Therefore, including gss.h in your project is recommended over gss/api.h. If using gss.h instead of gss/api.h causes problems, it should be regarded a bug.

You must include either file in all programs using the library, either directly or through some other header file, like this:

     #include <gss.h>

The name space of GSS is gss_* for function names, gss_* for data types and GSS_* for other symbols. In addition the same name prefixes with one prepended underscore are reserved for internal use and should never be used by an application.

Each supported GSS mechanism may want to expose mechanism specific functionality, and can do so through one or more header files under the gss/ directory. The Kerberos 5 mechanism uses the file gss/krb5.h, but again, it is included (with C++ namespace fixes)

from gss.h.

2.2 Initialization

GSS does not need to be initialized before it can be used.

In order to take advantage of the internationalisation features in GSS, e.g. translated error messages, the application must set the current locale using setlocale() before calling, e.g., gss_display_status(). This is typically done in main() as in the following example.

     #include <gss.h>
     #include <locale.h>
     ...
       setlocale (LC_ALL, "");

2.3 Version Check

It is often desirable to check that the version of GSS used is indeed one which fits all requirements. Even with binary compatibility new features may have been introduced but due to problem with the dynamic linker an old version is actually used. So you may want to check that the version is okay right after program startup. The function is called gss_ check_version() and is described formally in See Chapter 4 [Extended GSS API], page 57.

The normal way to use the function is to put something similar to the following early in

your main():
     #include <gss.h>
     ...
       if (!gss_check_version (GSS_VERSION))
         {
           printf ("gss_check_version() failed:\n"
                   "Header file incompatible with shared library.\n");
           exit(EXIT_FAILURE);
         }

2.4 Building the source

If you want to compile a source file that includes the gss.h header file, you must make sure that the compiler can find it in the directory hierarchy. This is accomplished by adding the path to the directory in which the header file is located to the compilers include file search path (via the -I option).

However, the path to the include file is determined at the time the source is configured. To solve this problem, GSS uses the external package pkg-config that knows the path to the include file and other configuration options. The options that need to be added to the compiler invocation at compile time are output by the --cflags option to pkg-config gss. The following example shows how it can be used at the command line:

     gcc -c foo.c `pkg-config gss --cflags`

Adding the output of ‘pkg-config gss --cflags’ to the compilers command line will ensure that the compiler can find the gss.h header file.

A similar problem occurs when linking the program with the library. Again, the compiler has to find the library files. For this to work, the path to the library files has to be added to the library search path (via the -L option). For this, the option --libs to pkg-config gss can be used. For convenience, this option also outputs all other options that are required to link the program with the GSS libarary (for instance, the ‘-lshishi’ option). The example shows how to link foo.o with GSS into a program foo.

     gcc -o foo foo.o `pkg-config gss --libs`

Of course you can also combine both examples to a single command by specifying both

options to pkg-config:
     gcc -o foo foo.c `pkg-config gss --cflags --libs`

2.5 Out of Memory handling

The GSS API does not have a standard error code for the out of memory error condition. This library will return GSS_S_FAILURE and set minor_status to ENOMEM.

3 Standard GSS API

3.1 Simple Data Types

The following conventions are used by the GSS-API C-language bindings:

3.1.1 Integer types

GSS-API uses the following integer data type:

   OM_uint32    32-bit unsigned integer

3.1.2 String and similar data

Many of the GSS-API routines take arguments and return values that describe contiguous octet-strings. All such data is passed between the GSS-API and the caller using the gss_ buffer_t data type. This data type is a pointer to a buffer descriptor, which consists of a length field that contains the total number of bytes in the datum, and a value field which contains a pointer to the actual datum:

   typedef struct gss_buffer_desc_struct {
      size_t    length;
      void      *value;
   } gss_buffer_desc, *gss_buffer_t;

Storage for data returned to the application by a GSS-API routine using the gss_ buffer_t conventions is allocated by the GSS-API routine. The application may free this storage by invoking the gss_release_buffer routine. Allocation of the gss_buffer_desc object is always the responsibility of the application; unused gss_buffer_desc objects may be initialized to the value GSS_C_EMPTY_BUFFER.

3.1.2.1 Opaque data types

Certain multiple-word data items are considered opaque data types at the GSS-API, because their internal structure has no significance either to the GSS-API or to the caller. Examples of such opaque data types are the input token parameter to gss_init_sec_context (which is opaque to the caller), and the input message parameter to gss_wrap (which is opaque to the GSS-API). Opaque data is passed between the GSS-API and the application using the

gss_buffer_t datatype.

3.1.2.2 Character strings

Certain multiple-word data items may be regarded as simple ISO Latin-1 character strings. Examples are the printable strings passed to gss_import_name via the input name buffer parameter. Some GSS-API routines also return character strings. All such character strings are passed between the application and the GSS-API implementation using the gss_buffer_t datatype, which is a pointer to a gss_buffer_desc object.

When a gss_buffer_desc object describes a printable string, the length field of the gss_buffer_desc should only count printable characters within the string. In particular, a trailing NUL character should NOT be included in the length count, nor should either the GSS-API implementation or the application assume the presence of an uncounted trailing NUL.

3.1.3 Object Identifiers

Certain GSS-API procedures take parameters of the type gss_OID, or Object identifier. This is a type containing ISO-defined tree- structured values, and is used by the GSS-API caller to select an underlying security mechanism and to specify namespaces. A value of type gss_OID has the following structure:

   typedef struct gss_OID_desc_struct {
      OM_uint32   length;
      void        *elements;
   } gss_OID_desc, *gss_OID;

The elements field of this structure points to the first byte of an octet string containing the ASN.1 BER encoding of the value portion of the normal BER TLV encoding of the gss_OID. The length field contains the number of bytes in this value. For example, the gss_

OID value corresponding to iso(1) identified-organization(3) icd-ecma(12) member-
company(2) dec(1011) cryptoAlgorithms(7) DASS(5), meaning the DASS X.509 authen-

tication mechanism, has a length field of 7 and an elements field pointing to seven octets containing the following octal values: 53,14,2,207,163,7,5. GSS-API implementations should provide constant gss_OID values to allow applications to request any supported mechanism, although applications are encouraged on portability grounds to accept the default mechanism. gss_OID values should also be provided to allow applications to specify particular name types (see section 3.10). Applications should treat gss_OID_desc values returned by GSS-API routines as read-only. In particular, the application should not attempt to deallocate them with free().

3.1.4 Object Identifier Sets

Certain GSS-API procedures take parameters of the type gss_OID_set. This type represents one or more object identifiers (see [Object Identifiers], page 10). A gss_OID_set object has the following structure:

   typedef struct gss_OID_set_desc_struct {
      size_t    count;
      gss_OID   elements;
   } gss_OID_set_desc, *gss_OID_set;

The count field contains the number of OIDs within the set. The elements field is a pointer to an array of gss_OID_desc objects, each of which describes a single OID. gss_ OID_set values are used to name the available mechanisms supported by the GSS-API, to request the use of specific mechanisms, and to indicate which mechanisms a given credential supports.

All OID sets returned to the application by GSS-API are dynamic objects (the gss_ OID_set_desc, the "elements" array of the set, and the "elements" array of each member OID are all dynamically allocated), and this storage must be deallocated by the application

using the gss_release_oid_set routine.

3.2 Complex Data Types

3.2.1 Credentials

A credential handle is a caller-opaque atomic datum that identifies a GSS-API credential data structure. It is represented by the caller- opaque type gss_cred_id_t.

GSS-API credentials can contain mechanism-specific principal authentication data for multiple mechanisms. A GSS-API credential is composed of a set of credential-elements, each of which is applicable to a single mechanism. A credential may contain at most one credential-element for each supported mechanism. A credential-element identifies the data needed by a single mechanism to authenticate a single principal, and conceptually contains two credential-references that describe the actual mechanism-specific authentication data, one to be used by GSS-API for initiating contexts, and one to be used for accepting contexts. For mechanisms that do not distinguish between acceptor and initiator credentials, both references would point to the same underlying mechanism-specific authentication data.

Credentials describe a set of mechanism-specific principals, and give their holder the ability to act as any of those principals. All principal identities asserted by a single GSS-API credential should belong to the same entity, although enforcement of this property is an implementation-specific matter. The GSS-API does not make the actual credentials available to applications; instead a credential handle is used to identify a particular credential, held internally by GSS-API. The combination of GSS-API credential handle and mechanism identifies the principal whose identity will be asserted by the credential when used with that mechanism.

   The gss_init_sec_context and gss_accept_sec_context routines allow the value

GSS_C_NO_CREDENTIAL to be specified as their credential handle parameter. This special credential-handle indicates a desire by the application to act as a default principal.

3.2.2 Contexts

The gss_ctx_id_t data type contains a caller-opaque atomic value that identifies one end of a GSS-API security context.

The security context holds state information about each end of a peer communication, including cryptographic state information.

3.2.3 Authentication tokens

A token is a caller-opaque type that GSS-API uses to maintain synchronization between the context data structures at each end of a GSS-API security context. The token is a cryptographically protected octet-string, generated by the underlying mechanism at one end of a GSS-API security context for use by the peer mechanism at the other end. Encapsulation (if required) and transfer of the token are the responsibility of the peer applications. A token is passed between the GSS-API and the application using the gss_buffer_t conventions.

3.2.4 Interprocess tokens

Certain GSS-API routines are intended to transfer data between processes in multi-process programs. These routines use a caller-opaque octet-string, generated by the GSS-API in one process for use by the GSS-API in another process. The calling application is responsible for transferring such tokens between processes in an OS-specific manner. Note that, while GSS-API implementors are encouraged to avoid placing sensitive information within interprocess tokens, or to cryptographically protect them, many implementations will be unable to avoid placing key material or other sensitive data within them. It is the application’s responsibility to ensure that interprocess tokens are protected in transit, and transferred only to processes that are trustworthy. An interprocess token is passed between the GSS-API and the application using the gss_buffer_t conventions.

3.2.5 Names

A name is used to identify a person or entity. GSS-API authenticates the relationship between a name and the entity claiming the name.

Since different authentication mechanisms may employ different namespaces for identifying their principals, GSSAPI’s naming support is necessarily complex in multi-mechanism environments (or even in some single-mechanism environments where the underlying mechanism supports multiple namespaces).

Two distinct representations are defined for names:

• An internal form. This is the GSS-API "native" format for names, represented by the implementation-specific gss_name_t type. It is opaque to GSS-API callers. A single gss_name_t object may contain multiple names from different namespaces, but all names should refer to the same entity. An example of such an internal name would be the name returned from a call to the gss_inquire_cred routine, when applied to a credential containing credential elements for multiple authentication mechanisms employing different namespaces. This gss_name_t object will contain a distinct name for the entity for each authentication mechanism. For GSS-API implementations supporting multiple namespaces, objects of type gss_ name_t must contain sufficient information to determine the namespace to which each primitive name belongs.

• Mechanism-specific contiguous octet-string forms. A format capable of containing a single name (from a single namespace). Contiguous string names are always accompanied by an object identifier specifying the namespace to which the name belongs, and their format is dependent on the authentication mechanism that employs the name. Many, but not all, contiguous string names will be printable, and may therefore be used by GSS-API applications for communication with their users.

Routines (gss_import_name and gss_display_name) are provided to convert names between contiguous string representations and the internal gss_name_t type. gss_import_ name may support multiple syntaxes for each supported namespace, allowing users the freedom to choose a preferred name representation. gss_display_name should use an implementation-chosen printable syntax for each supported name-type.

If an application calls gss_display_name, passing the internal name resulting from a call to gss_import_name, there is no guarantee the resulting contiguous string name will be the same as the original imported string name. Nor do name-space identifiers necessarily survive unchanged after a journey through the internal name-form. An example of this might be a mechanism that authenticates X.500 names, but provides an algorithmic mapping of Internet DNS names into X.500. That mechanism’s implementation of gss_import_name might, when presented with a DNS name, generate an internal name that contained both the original DNS name and the equivalent X.500 name. Alternatively, it might only store the X.500 name. In the latter case, gss_display_name would most likely generate a printable X.500 name, rather than the original DNS name.

The process of authentication delivers to the context acceptor an internal name. Since this name has been authenticated by a single mechanism, it contains only a single name (even if the internal name presented by the context initiator to gss_init_sec_context had multiple components). Such names are termed internal mechanism names, or "MN"s and the names emitted by gss_accept_sec_context are always of this type. Since some applications may require MNs without wanting to incur the overhead of an authentication operation, a second function, gss_canonicalize_name, is provided to convert a general internal name into an MN.

Comparison of internal-form names may be accomplished via the gss_compare_name routine, which returns true if the two names being compared refer to the same entity. This removes the need for the application program to understand the syntaxes of the various printable names that a given GSS-API implementation may support. Since GSS-API assumes that all primitive names contained within a given internal name refer to the same entity, gss_compare_name can return true if the two names have at least one primitive name in common. If the implementation embodies knowledge of equivalence relationships between names taken from different namespaces, this knowledge may also allow successful comparison of internal names containing no overlapping primitive elements.

When used in large access control lists, the overhead of invoking gss_import_name and gss_compare_name on each name from the ACL may be prohibitive. As an alternative way of supporting this case, GSS-API defines a special form of the contiguous string name which may be compared directly (e.g. with memcmp()). Contiguous names suitable for comparison are generated by the gss_export_name routine, which requires an MN as input. Exported names may be re- imported by the gss_import_name routine, and the resulting internal name will also be an MN. The gss_OID constant GSS_C_NT_EXPORT_NAME indentifies the "export name" type, and the value of this constant is given in Appendix A. Structurally, an exported name object consists of a header containing an OID identifying the mechanism that authenticated the name, and a trailer containing the name itself, where the syntax of the trailer is defined by the individual mechanism specification. The precise format of an export name is defined in the language-independent GSS-API specification [GSSAPI].

Note that the results obtained by using gss_compare_name will in general be different from those obtained by invoking gss_canonicalize_name and gss_export_name, and then comparing the exported names. The first series of operation determines whether two (unauthenticated) names identify the same principal; the second whether a particular mechanism would authenticate them as the same principal. These two operations will in general give the same results only for MNs.

The gss_name_t datatype should be implemented as a pointer type. To allow the compiler to aid the application programmer by performing type-checking, the use of (void *) is discouraged. A pointer to an implementation-defined type is the preferred choice.

Storage is allocated by routines that return gss_name_t values. A procedure, gss_ release_name, is provided to free storage associated with an internal-form name.

3.2.6 Channel Bindings

GSS-API supports the use of user-specified tags to identify a given context to the peer application. These tags are intended to be used to identify the particular communications channel that carries the context. Channel bindings are communicated to the GSS-API using the following structure:

typedef struct gss_channel_bindings_struct {
   OM_uint32       initiator_addrtype;
   gss_buffer_desc initiator_address;
   OM_uint32       acceptor_addrtype;
   gss_buffer_desc acceptor_address;
   gss_buffer_desc application_data;
} *gss_channel_bindings_t;

The initiator addrtype and acceptor addrtype fields denote the type of addresses contained in the initiator address and acceptor address buffers. The address type should be one of the following:

GSS_C_AF_UNSPEC     Unspecified address type
GSS_C_AF_LOCAL      Host-local address type
GSS_C_AF_INET       Internet address type (e.g. IP)
GSS_C_AF_IMPLINK    ARPAnet IMP address type
GSS_C_AF_PUP        pup protocols (eg BSP) address type
GSS_C_AF_CHAOS      MIT CHAOS protocol address type
GSS_C_AF_NS         XEROX NS address type
GSS_C_AF_NBS        nbs address type
GSS_C_AF_ECMA       ECMA address type
GSS_C_AF_DATAKIT    datakit protocols address type
GSS_C_AF_CCITT      CCITT protocols
GSS_C_AF_SNA        IBM SNA address type
GSS_C_AF_DECnet     DECnet address type
GSS_C_AF_DLI        Direct data link interface address type
GSS_C_AF_LAT        LAT address type
GSS_C_AF_HYLINK     NSC Hyperchannel address type
GSS_C_AF_APPLETALK  AppleTalk address type
GSS_C_AF_BSC        BISYNC 2780/3780 address type
GSS_C_AF_DSS        Distributed system services address type
GSS_C_AF_OSI        OSI TP4 address type
GSS_C_AF_X25        X.25
GSS_C_AF_NULLADDR   No address specified

Note that these symbols name address families rather than specific addressing formats. For address families that contain several alternative address forms, the initiator address and acceptor address fields must contain sufficient information to determine which address form is used. When not otherwise specified, addresses should be specified in network byte-order (that is, native byte-ordering for the address family).

Conceptually, the GSS-API concatenates the initiator addrtype, initiator address, acceptor addrtype, acceptor address and application data to form an octet string. The mechanism calculates a MIC over this octet string, and binds the MIC to the context establishment token emitted by gss_init_sec_context. The same bindings are presented by the context acceptor to gss_accept_sec_context, and a MIC is calculated in the same way. The calculated MIC is compared with that found in the token, and if the MICs differ,

gss_accept_sec_context will return a GSS_S_BAD_BINDINGS error, and the context will

not be established. Some mechanisms may include the actual channel binding data in the token (rather than just a MIC); applications should therefore not use confidential data as channel-binding components.

Individual mechanisms may impose additional constraints on addresses and address types that may appear in channel bindings. For example, a mechanism may verify that the initiator address field of the channel bindings presented to gss_init_sec_context contains the correct network address of the host system. Portable applications should therefore ensure that they either provide correct information for the address fields, or omit addressing information, specifying GSS_C_AF_NULLADDR as the address-types.

3.3 Optional Parameters

Various parameters are described as optional. This means that they follow a convention whereby a default value may be requested. The following conventions are used for omitted parameters. These conventions apply only to those parameters that are explicitly documented as optional.

• gss buffer t types. Specify GSS C NO BUFFER as a value. For an input parameter this signifies that default behavior is requested, while for an output parameter it indicates that the information that would be returned via the parameter is not required by the application.

• Integer types (input). Individual parameter documentation lists values to be used to indicate default actions.

• Integer types (output). Specify NULL as the value for the pointer.

• Pointer types. Specify NULL as the value.

• Object IDs. Specify GSS C NO OID as the value.

• Object ID Sets. Specify GSS C NO OID SET as the value.

• Channel Bindings. Specify GSS C NO CHANNEL BINDINGS to indicate that channel bindings are not to be used.

3.4 Error Handling

Every GSS-API routine returns two distinct values to report status information to the caller: GSS status codes and Mechanism status codes.

3.4.1 GSS status codes

GSS-API routines return GSS status codes as their OM_uint32 function value. These codes indicate errors that are independent of the underlying mechanism(s) used to provide the security service. The errors that can be indicated via a GSS status code are either generic API routine errors (errors that are defined in the GSS-API specification) or calling errors (errors that are specific to these language bindings).

A GSS status code can indicate a single fatal generic API error from the routine and a single calling error. In addition, supplementary status information may be indicated via the setting of bits in the supplementary info field of a GSS status code.

These errors are encoded into the 32-bit GSS status code as follows:

      MSB                                                        LSB
      |------------------------------------------------------------|

   |  Calling Error | Routine Error  |    Supplementary Info    |
   |------------------------------------------------------------|
Bit 31            24 23            16 15                       0

Hence if a GSS-API routine returns a GSS status code whose upper 16 bits contain a non-zero value, the call failed. If the calling error field is non-zero, the invoking application’s call of the routine was erroneous. Calling errors are defined in table 3-1. If the routine error field is non-zero, the routine failed for one of the routine- specific reasons listed below in table 3-2. Whether or not the upper 16 bits indicate a failure or a success, the routine may indicate additional information by setting bits in the supplementary info field of the status code. The meaning of individual bits is listed below in table 3-3.

Table 3-1  Calling Errors
Name                   Value in field           Meaning
----                   --------------           -------
GSS_S_CALL_INACCESSIBLE_READ  1       A required input parameter
                                      could not be read
GSS_S_CALL_INACCESSIBLE_WRITE 2       A required output parameter
                                       could not be written.
GSS_S_CALL_BAD_STRUCTURE      3       A parameter was malformed
Table 3-2  Routine Errors
Name                   Value in field           Meaning
----                   --------------           -------
GSS_S_BAD_MECH                1       An unsupported mechanism
                                      was requested
GSS_S_BAD_NAME                2       An invalid name was
                                      supplied
GSS_S_BAD_NAMETYPE            3       A supplied name was of an
                                      unsupported type
GSS_S_BAD_BINDINGS            4       Incorrect channel bindings
                                      were supplied
GSS_S_BAD_STATUS              5       An invalid status code was
                                      supplied
GSS_S_BAD_MIC GSS_S_BAD_SIG   6       A token had an invalid MIC
GSS_S_NO_CRED                 7       No credentials were
                                      supplied, or the
                                      credentials were
                                      unavailable or
                                      inaccessible.
GSS_S_NO_CONTEXT              8       No context has been
                                      established
GSS_S_DEFECTIVE_TOKEN         9       A token was invalid
GSS_S_DEFECTIVE_CREDENTIAL   10       A credential was invalid
GSS_S_CREDENTIALS_EXPIRED    11       The referenced credentials
                                      have expired
GSS_S_CONTEXT_EXPIRED        12       The context has expired

GSS_S_FAILURE                13       Miscellaneous failure (see
                                      text)
GSS_S_BAD_QOP                14       The quality-of-protection
                                      requested could not be
                                      provided
GSS_S_UNAUTHORIZED           15       The operation is forbidden
                                      by local security policy
GSS_S_UNAVAILABLE            16       The operation or option is
                                      unavailable
GSS_S_DUPLICATE_ELEMENT      17       The requested credential
                                      element already exists
GSS_S_NAME_NOT_MN            18       The provided name was not a
                                      mechanism name
Table 3-3  Supplementary Status Bits
Name                   Bit Number           Meaning
----                   ----------           -------
GSS_S_CONTINUE_NEEDED   0 (LSB)   Returned only by
                                  gss_init_sec_context or
                                  gss_accept_sec_context. The
                                  routine must be called again
                                  to complete its function.
                                  See routine documentation for
                                  detailed description
GSS_S_DUPLICATE_TOKEN   1         The token was a duplicate of
                                  an earlier token
GSS_S_OLD_TOKEN         2         The token's validity period
                                  has expired
GSS_S_UNSEQ_TOKEN       3         A later token has already been
                                  processed
GSS_S_GAP_TOKEN         4         An expected per-message token
                                  was not received

The routine documentation also uses the name GSS S COMPLETE, which is a zero value, to indicate an absence of any API errors or supplementary information bits.

All GSS S xxx symbols equate to complete OM_uint32 status codes, rather than to bitfield values. For example, the actual value of the symbol GSS_S_BAD_NAMETYPE (value 3 in the routine error field) is 3<<16. The macros GSS_CALLING_ERROR, GSS_ROUTINE_ ERROR and GSS_SUPPLEMENTARY_INFO are provided, each of which takes a GSS status code and removes all but the relevant field. For example, the value obtained by applying GSS_ ROUTINE_ERROR to a status code removes the calling errors and supplementary info fields, leaving only the routine errors field. The values delivered by these macros may be directly compared with a GSS_S_xxx symbol of the appropriate type. The macro GSS_ERROR is also provided, which when applied to a GSS status code returns a non-zero value if the status code indicated a calling or routine error, and a zero value otherwise. All macros defined by GSS-API evaluate their argument(s) exactly once.

A GSS-API implementation may choose to signal calling errors in a platform-specific manner instead of, or in addition to the routine value; routine errors and supplementary info should be returned via major status values only.

The GSS major status code GSS_S_FAILURE is used to indicate that the underlying mechanism detected an error for which no specific GSS status code is defined. The mechanismspecific status code will provide more details about the error.

In addition to the explicit major status codes for each API function, the code GSS_S_FAILURE may be returned by any routine, indicating an implementation-specific or mechanism-specific error condition, further details of which are reported via the

minor_status parameter.

3.4.2 Mechanism-specific status codes

GSS-API routines return a minor status parameter, which is used to indicate specialized errors from the underlying security mechanism. This parameter may contain a single mechanism-specific error, indicated by a OM_uint32 value.

The minor status parameter will always be set by a GSS-API routine, even if it returns a calling error or one of the generic API errors indicated above as fatal, although most other output parameters may remain unset in such cases. However, output parameters that are expected to return pointers to storage allocated by a routine must always be set by the routine, even in the event of an error, although in such cases the GSS-API routine may elect to set the returned parameter value to NULL to indicate that no storage was actually allocated. Any length field associated with such pointers (as in a gss_buffer_ desc structure) should also be set to zero in such cases.

3.5 Credential Management

   GSS-API Credential-management Routines
   Routine                         Function
   -------                         --------
   gss_acquire_cred                Assume a global identity; Obtain
                                   a GSS-API credential handle for
                                   pre-existing credentials.
   gss_add_cred                    Construct credentials
                                   incrementally.
   gss_inquire_cred                Obtain information about a
                                   credential.
   gss_inquire_cred_by_mech        Obtain per-mechanism information
                                   about a credential.
   gss_release_cred                Discard a credential handle.

gss acquire cred

[Function] OM_uint32 gss_acquire_cred (OM uint32 * minor_status, const

gss name t desired_name, OM uint32 time_req, const gss OID set desired_mechs, gss cred usage t cred_usage, gss cred id t *

   output_cred_handle, gss OID set * actual_mechs, OM uint32 *
   time_rec)

minor status: (integer, modify) Mechanism specific status code. desired name: (gss name t, read) Name of principal whose credential should be acquired. time req: (Integer, read, optional) Number of seconds that credentials should remain valid. Specify GSS C INDEFINITE to request that the credentials have the maximum permitted lifetime. desired mechs: (Set of Object IDs, read, optional) Set of underlying security mechanisms that may be used. GSS C NO OID SET may be used to obtain an implementation-specific default. cred usage: (gss cred usage t, read) GSS C BOTH - Credentials may be used either to initiate or accept security contexts. GSS C INITIATE - Credentials will only be used to initiate security contexts. GSS C ACCEPT - Credentials will only be used to accept security contexts. output cred handle: (gss cred id t, modify) The returned credential handle. Resources associated with this credential handle must be released by the application after use with a call to gss release cred(). actual mechs: (Set of Object IDs, modify, optional) The set of mechanisms for which the credential is valid. Storage associated with the returned OID-set must be released by the application after use with a call to gss release oid set(). Specify NULL if not required. time rec: (Integer, modify, optional) Actual number of seconds for which the returned credentials will remain valid. If the implementation does not support expiration of credentials, the value GSS C INDEFINITE will be returned. Specify NULL if not required. Allows an application to acquire a handle for a pre-existing credential by name. GSS-API implementations must impose a local access-control policy on callers of this routine to prevent unauthorized callers from acquiring credentials to which they are not entitled. This routine is not intended to provide a "login to the network" function, as such a function would involve the creation of new credentials rather than merely acquiring a handle to existing credentials. Such functions, if required, should be defined in implementation-specific extensions to the API. If desired name is GSS C NO NAME, the call is interpreted as a request for a credential handle that will invoke default behavior when passed to gss init sec context() (if cred usage is GSS C INITIATE or GSS C BOTH) or gss accept sec context() (if cred usage is GSS C ACCEPT or GSS C BOTH). Mechanisms should honor the desired mechs parameter, and return a credential that is suitable to use only with the requested mechanisms. An exception to this is the case where one underlying credential element can be shared by multiple mechanisms; in this case it is permissible for an implementation to indicate all mechanisms with which the credential element may be used. If desired mechs is an empty set, behavior is undefined. This routine is expected to be used primarily by context acceptors, since implementations are likely to provide mechanism-specific ways of obtaining GSS-API initiator credentials from the system login process. Some implementations may therefore not support the acquisition of GSS C INITIATE or GSS C BOTH credentials via gss acquire cred for any name other than GSS C NO NAME, or a name produced by applying either gss inquire cred to a valid credential, or gss inquire context to an active context. If credential acquisition is time-consuming for a mechanism, the mechanism may choose to delay the actual acquisition until the credential is required (e.g. by gss init sec context or gss accept sec context). Such mechanism-specific implementation decisions should be invisible to the calling application; thus a call of gss inquire cred immediately following the call of gss acquire cred must return valid credential data, and may therefore incur the overhead of a deferred credential acquisition. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_BAD_MECH: Unavailable mechanism requested. GSS_S_BAD_NAMETYPE: Type contained within desired name parameter is not supported. GSS_S_BAD_NAME: Value supplied for desired name parameter is ill formed. GSS_S_CREDENTIALS_EXPIRED: The credentials could not be acquired Because they have expired. GSS_S_NO_CRED: No credentials were found for the specified name.

gss add cred

[Function] OM_uint32 gss_add_cred (OM uint32 * minor_status, const

gss cred id t input_cred_handle, const gss name t desired_name, const gss OID desired_mech, gss cred usage t cred_usage, OM uint32

   initiator_time_req, OM uint32 acceptor_time_req, gss cred id t *
   output_cred_handle, gss OID set * actual_mechs, OM uint32 *
   initiator_time_rec, OM uint32 * acceptor_time_rec)

minor status: (integer, modify) Mechanism specific status code. input cred handle: (gss cred id t, read, optional) The credential to which a credential-element will be added. If GSS C NO CREDENTIAL is specified, the routine will compose the new credential based on default behavior (see text). Note that, while the credential-handle is not modified by gss add cred(), the underlying credential will be modified if output credential handle is NULL. desired name: (gss name t, read.) Name of principal whose credential should be acquired. desired mech: (Object ID, read) Underlying security mechanism with which the credential may be used. cred usage: (gss cred usage t, read) GSS C BOTH - Credential may be used either to initiate or accept security contexts. GSS C INITIATE - Credential will only be used to initiate security contexts. GSS C ACCEPT - Credential will only be used to accept security contexts. initiator time req: (Integer, read, optional) number of seconds that the credential should remain valid for initiating security contexts. This argument is ignored if the composed credentials are of type GSS C ACCEPT. Specify GSS C INDEFINITE to request that the credentials have the maximum permitted initiator lifetime. acceptor time req: (Integer, read, optional) number of seconds that the credential should remain valid for accepting security contexts. This argument is ignored if the composed credentials are of type GSS C INITIATE. Specify GSS C INDEFINITE to request that the credentials have the maximum permitted initiator lifetime. output cred handle: (gss cred id t, modify, optional) The returned credential handle, containing the new credential-element and all the credential-elements from input cred handle. If a valid pointer to a gss cred id t is supplied for this parameter, gss add cred creates a new credential handle containing all credential-elements from the input cred handle and the newly acquired credential-element; if NULL is specified for this parameter, the newly acquired credential-element will be added to the credential identified by input cred handle. The resources associated with any credential handle returned via this parameter must be released by the application after use with a call to gss release cred(). actual mechs: (Set of Object IDs, modify, optional) The complete set of mechanisms for which the new credential is valid. Storage for the returned OID-set must be freed by the application after use with a call to gss release oid set(). Specify NULL if not required. initiator time rec: (Integer, modify, optional) Actual number of seconds for which the returned credentials will remain valid for initiating contexts using the specified mechanism. If the implementation or mechanism does not support expiration of credentials, the value GSS C INDEFINITE will be returned. Specify NULL if not required acceptor time rec: (Integer, modify, optional) Actual number of seconds for which the returned credentials will remain valid for accepting security contexts using the specified mechanism. If the implementation or mechanism does not support expiration of credentials, the value GSS C INDEFINITE will be returned. Specify NULL if not required Adds a credential-element to a credential. The credential-element is identified by the name of the principal to which it refers. GSS-API implementations must impose a local access-control policy on callers of this routine to prevent unauthorized callers from acquiring credential-elements to which they are not entitled. This routine is not intended to provide a "login to the network" function, as such a function would involve the creation of new mechanism-specific authentication data, rather than merely acquiring a GSS-API handle to existing data. Such functions, if required, should be defined in implementation-specific extensions to the API. If desired name is GSS C NO NAME, the call is interpreted as a request to add a credential element that will invoke default behavior when passed to gss init sec context()

(if cred usage is GSS C INITIATE or GSS C BOTH) or gss accept sec context() (if cred usage is GSS C ACCEPT or GSS C BOTH).

This routine is expected to be used primarily by context acceptors, since implementations are likely to provide mechanism-specific ways of obtaining GSS-API initiator credentials from the system login process. Some implementations may therefore not support the acquisition of GSS C INITIATE or GSS C BOTH credentials via gss acquire cred for any name other than GSS C NO NAME, or a name produced by applying either gss inquire cred to a valid credential, or gss inquire context to an active context.

If credential acquisition is time-consuming for a mechanism, the mechanism may choose to delay the actual acquisition until the credential is required (e.g. by gss init sec context or gss accept sec context). Such mechanism-specific implementation decisions should be invisible to the calling application; thus a call of gss inquire cred immediately following the call of gss add cred must return valid credential data, and may therefore incur the overhead of a deferred credential acquisition.

This routine can be used to either compose a new credential containing all credentialelements of the original in addition to the newly-acquire credential-element, or to add the new credential- element to an existing credential. If NULL is specified for the output cred handle parameter argument, the new credential-element will be added to the credential identified by input cred handle; if a valid pointer is specified for the output cred handle parameter, a new credential handle will be created.

If GSS C NO CREDENTIAL is specified as the input cred handle, gss add cred will compose a credential (and set the output cred handle parameter accordingly) based on default behavior. That is, the call will have the same effect as if the application had first made a call to gss acquire cred(), specifying the same usage and passing GSS C NO NAME as the desired name parameter to obtain an explicit credential handle embodying default behavior, passed this credential handle to gss add cred(), and finally called gss release cred() on the first credential handle.

If GSS C NO CREDENTIAL is specified as the input cred handle parameter, a non-NULL output cred handle must be supplied.

Return value:

GSS_S_COMPLETE: Successful completion.

GSS_S_BAD_MECH: Unavailable mechanism requested.

GSS_S_BAD_NAMETYPE: Type contained within desired name parameter is not supported.

GSS_S_BAD_NAME: Value supplied for desired name parameter is ill-formed.

GSS_S_DUPLICATE_ELEMENT: The credential already contains an element for the requested mechanism with overlapping usage and validity period.

GSS_S_CREDENTIALS_EXPIRED: The required credentials could not be added because they have expired.

GSS_S_NO_CRED: No credentials were found for the specified name.

gss inquire cred

[Function] OM_uint32 gss_inquire_cred (OM uint32 * minor_status, const

gss cred id t cred_handle, gss name t * name, OM uint32 * lifetime, gss cred usage t * cred_usage, gss OID set * mechanisms)

minor status: (integer, modify) Mechanism specific status code.

cred handle: (gss cred id t, read) A handle that refers to the target credential. Specify GSS C NO CREDENTIAL to inquire about the default initiator principal.

name: (gss name t, modify, optional) The name whose identity the credential asserts. Storage associated with this name should be freed by the application after use with a call to gss release name(). Specify NULL if not required.

lifetime: (Integer, modify, optional) The number of seconds for which the credential will remain valid. If the credential has expired, this parameter will be set to zero. If the implementation does not support credential expiration, the value GSS C INDEFINITE will be returned. Specify NULL if not required.

cred usage: (gss cred usage t, modify, optional) How the credential may be used. One of the following: GSS C INITIATE, GSS C ACCEPT, GSS C BOTH. Specify NULL if not required.

mechanisms: (gss OID set, modify, optional) Set of mechanisms supported by the credential. Storage associated with this OID set must be freed by the application after use with a call to gss release oid set(). Specify NULL if not required.

Obtains information about a credential.

Return value:

GSS_S_COMPLETE: Successful completion.

GSS_S_NO_CRED: The referenced credentials could not be accessed.

GSS_S_DEFECTIVE_CREDENTIAL: The referenced credentials were invalid.

GSS_S_CREDENTIALS_EXPIRED: The referenced credentials have expired. If the lifetime parameter was not passed as NULL, it will be set to 0.

gss inquire cred by mech

[Function] OM_uint32 gss_inquire_cred_by_mech (OM uint32 *
   minor_status, const gss cred id t cred_handle, const gss OID mech_type,

gss name t * name, OM uint32 * initiator_lifetime, OM uint32 *

   acceptor_lifetime, gss cred usage t * cred_usage)

minor status: (Integer, modify) Mechanism specific status code.

cred handle: (gss cred id t, read) A handle that refers to the target credential. Specify GSS C NO CREDENTIAL to inquire about the default initiator principal.

mech type: (gss OID, read) The mechanism for which information should be returned.

name: (gss name t, modify, optional) The name whose identity the credential asserts. Storage associated with this name must be freed by the application after use with a call to gss release name(). Specify NULL if not required. initiator lifetime: (Integer, modify, optional) The number of seconds for which the credential will remain capable of initiating security contexts under the specified mechanism. If the credential can no longer be used to initiate contexts, or if the credential usage for this mechanism is GSS C ACCEPT, this parameter will be set to zero. If the implementation does not support expiration of initiator credentials, the value GSS C INDEFINITE will be returned. Specify NULL if not required. acceptor lifetime: (Integer, modify, optional) The number of seconds for which the credential will remain capable of accepting security contexts under the specified mechanism. If the credential can no longer be used to accept contexts, or if the credential usage for this mechanism is GSS C INITIATE, this parameter will be set to zero. If the implementation does not support expiration of acceptor credentials, the value GSS C INDEFINITE will be returned. Specify NULL if not required. cred usage: (gss cred usage t, modify, optional) How the credential may be used with the specified mechanism. One of the following: GSS C INITIATE, GSS C ACCEPT, GSS C BOTH. Specify NULL if not required. Obtains per-mechanism information about a credential. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_NO_CRED: The referenced credentials could not be accessed. GSS_S_DEFECTIVE_CREDENTIAL: The referenced credentials were invalid. GSS_S_CREDENTIALS_EXPIRED: The referenced credentials have expired. If the lifetime parameter was not passed as NULL, it will be set to 0.

gss release cred

[Function] OM_uint32 gss_release_cred (OM uint32 * minor_status,

gss cred id t * cred_handle) minor status: (Integer, modify) Mechanism specific status code. cred handle: (gss cred id t, modify, optional) Opaque handle identifying credential to be released. If GSS C NO CREDENTIAL is supplied, the routine will complete successfully, but will do nothing. Informs GSS-API that the specified credential handle is no longer required by the application, and frees associated resources. The cred handle is set to GSS C NO CREDENTIAL on successful completion of this call. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_NO_CRED: Credentials could not be accessed.

3.6 Context-Level Routines

GSS-API Context-Level Routines
Routine                         Function
-------                         --------
gss_init_sec_context            Initiate a security context with

                              a peer application.
gss_accept_sec_context          Accept a security context
                              initiated by a peer application.
gss_delete_sec_context          Discard a security context.
gss_process_context_token       Process a token on a security
                              context from a peer application.
gss_context_time                Determine for how long a context
                              will remain valid.
gss_inquire_context             Obtain information about a
                              security context.
gss_wrap_size_limit             Determine token-size limit for
                              gss_wrap on a context.
gss_export_sec_context          Transfer a security context to
                              another process.
gss_import_sec_context          Import a transferred context.

gss init sec context

[Function] OM_uint32 gss_init_sec_context (OM uint32 * minor_status,

const gss cred id t initiator_cred_handle, gss ctx id t * context_handle, const gss name t target_name, const gss OID

   mech_type, OM uint32 req_flags, OM uint32 time_req, const

gss channel bindings t input_chan_bindings, const gss buffer t

   input_token, gss OID * actual_mech_type, gss buffer t output_token,

OM uint32 * ret_flags, OM uint32 * time_rec) minor status: (integer, modify) Mechanism specific status code. initiator cred handle: (gss cred id t, read, optional) Handle for credentials claimed. Supply GSS C NO CREDENTIAL to act as a default initiator principal. If no default initiator is defined, the function will return GSS S NO CRED. context handle: (gss ctx id t, read/modify) Context handle for new context. Supply GSS C NO CONTEXT for first call; use value returned by first call in continuation calls. Resources associated with this context-handle must be released by the application after use with a call to gss delete sec context(). target name: (gss name t, read) Name of target. mech type: (OID, read, optional) Object ID of desired mechanism. Supply GSS C NO OID to obtain an implementation specific default. req flags: (bit-mask, read) Contains various independent flags, each of which requests that the context support a specific service option. Symbolic names are provided for each flag, and the symbolic names corresponding to the required flags should be logically-ORed together to form the bit-mask value. See below for the flags. time req: (Integer, read, optional) Desired number of seconds for which context should remain valid. Supply 0 to request a default validity period. input chan bindings: (channel bindings, read, optional) Application-specified bindings. Allows application to securely bind channel identification information to the security context. Specify GSS C NO CHANNEL BINDINGS if channel bindings are not used. input token: (buffer, opaque, read, optional) Token received from peer application. Supply GSS C NO BUFFER, or a pointer to a buffer containing the value GSS C EMPTY BUFFER on initial call. actual mech type: (OID, modify, optional) Actual mechanism used. The OID returned via this parameter will be a pointer to static storage that should be treated as read-only; In particular the application should not attempt to free it. Specify NULL if not required. output token: (buffer, opaque, modify) Token to be sent to peer application. If the length field of the returned buffer is zero, no token need be sent to the peer application. Storage associated with this buffer must be freed by the application after use with a call to gss release buffer(). ret flags: (bit-mask, modify, optional) Contains various independent flags, each of which indicates that the context supports a specific service option. Specify NULL if not required. Symbolic names are provided for each flag, and the symbolic names corresponding to the required flags should be logically-ANDed with the ret flags value to test whether a given option is supported by the context. See below for the flags. time rec: (Integer, modify, optional) Number of seconds for which the context will remain valid. If the implementation does not support context expiration, the value GSS C INDEFINITE will be returned. Specify NULL if not required. Initiates the establishment of a security context between the application and a remote peer. Initially, the input token parameter should be specified either as GSS C NO BUFFER, or as a pointer to a gss buffer desc object whose length field contains the value zero. The routine may return a output token which should be transferred to the peer application, where the peer application will present it to gss accept sec context. If no token need be sent, gss init sec context will indicate this by setting the length field of the output token argument to zero. To complete the context establishment, one or more reply tokens may be required from the peer application; if so, gss init sec context will return a status containing the supplementary information bit GSS S CONTINUE NEEDED. In this case, gss init sec context should be called again when the reply token is received from the peer application, passing the reply token to gss init sec context via the input token parameters. Portable applications should be constructed to use the token length and return status to determine whether a token needs to be sent or waited for. Thus a typical portable caller should always invoke gss init sec context within a loop:

     int context_established = 0;
     gss_ctx_id_t context_hdl = GSS_C_NO_CONTEXT;
            ...
     input_token->length = 0;
     while (!context_established) {
       maj_stat = gss_init_sec_context(&min_stat,
                                       cred_hdl,
                                       &context_hdl,
                                       target_name,

desired_mech,
desired_services,
desired_time,
input_bindings,
input_token,
&actual_mech,
output_token,
&actual_services,
&actual_time);
if (GSS_ERROR(maj_stat)) {
report_error(maj_stat, min_stat);
};
if (output_token->length != 0) {
send_token_to_peer(output_token);
gss_release_buffer(&min_stat, output_token)
};
if (GSS_ERROR(maj_stat)) {
if (context_hdl != GSS_C_NO_CONTEXT)
gss_delete_sec_context(&min_stat,
&context_hdl,
GSS_C_NO_BUFFER);
break;
};
if (maj_stat & GSS_S_CONTINUE_NEEDED) {
receive_token_from_peer(input_token);
} else {
context_established = 1;
};
};

Whenever the routine returns a major status that includes the value GSS S CONTINUE NEEDED, the context is not fully established and the following restrictions apply to the output parameters:

• The value returned via the time rec parameter is undefined unless the accompanying ret flags parameter contains the bit GSS C PROT READY FLAG, indicating that per-message services may be applied in advance of a successful completion status, the value returned via the actual mech type parameter is undefined until the routine returns a major status value of GSS S COMPLETE.

• The values of the GSS C DELEG FLAG, GSS C MUTUAL FLAG, GSS C REPLAY FLAG, GSS C SEQUENCE FLAG, GSS C CONF FLAG, GSS C INTEG FLAG and GSS C ANON FLAG bits returned via the ret flags parameter should contain the values that the implementation expects would be valid if context establishment were to succeed. In particular, if the application has requested a service such as delegation or anonymous authentication via the req flags argument, and such a service is unavailable from the underlying mechanism, gss init sec context should generate a token that will not provide the service, and indicate via the ret flags argument that the service will not be supported. The application may choose to abort the context establishment by calling gss delete sec context (if it cannot continue in the absence of the service), or it may choose to transmit the token and continue context establishment (if the service was merely desired but not mandatory).

• The values of the GSS C PROT READY FLAG and GSS C TRANS FLAG bits within ret flags should indicate the actual state at the time gss init sec context returns, whether or not the context is fully established.

• GSS-API implementations that support per-message protection are encouraged to set the GSS C PROT READY FLAG in the final ret flags returned to a caller (i.e. when accompanied by a GSS S COMPLETE status code). However, applications should not rely on this behavior as the flag was not defined in Version 1 of the GSS-API. Instead, applications should determine what per-message services are available after a successful context establishment according to the GSS C INTEG FLAG and GSS C CONF FLAG values.

• All other bits within the ret flags argument should be set to zero.

If the initial call of gss init sec context() fails, the implementation should not create a context object, and should leave the value of the context handle parameter set to GSS C NO CONTEXT to indicate this. In the event of a failure on a subsequent call, the implementation is permitted to delete the "half-built" security context (in which case it should set the context handle parameter to GSS C NO CONTEXT), but the preferred behavior is to leave the security context untouched for the application to delete (using gss delete sec context). During context establishment, the informational status bits GSS S OLD TOKEN and GSS S DUPLICATE TOKEN indicate fatal errors, and GSS-API mechanisms should always return them in association with a routine error of GSS S FAILURE. This requirement for pairing did not exist in version 1 of the GSS-API specification, so applications that wish to run over version 1 implementations must special-case these codes. The req_flags values:

      GSS_C_DELEG_FLAG

• True - Delegate credentials to remote peer.

• False - Don’t delegate.

      GSS_C_MUTUAL_FLAG

• True - Request that remote peer authenticate itself.

• False - Authenticate self to remote peer only.

      GSS_C_REPLAY_FLAG

• True - Enable replay detection for messages protected with gss wrap or gss get mic.

• False - Don’t attempt to detect replayed messages.

GSS_C_SEQUENCE_FLAG

• True - Enable detection of out-of-sequence protected messages.

• False - Don’t attempt to detect out-of-sequence messages.

GSS_C_CONF_FLAG

• True - Request that confidentiality service be made available (via gss wrap).

• False - No per-message confidentiality service is required.

GSS_C_INTEG_FLAG

• True - Request that integrity service be made available (via gss wrap or gss get mic).

• False - No per-message integrity service is required.

GSS_C_ANON_FLAG

• True - Do not reveal the initiator’s identity to the acceptor.

• False - Authenticate normally.

The ret_flags values:

GSS_C_DELEG_FLAG

• True - Credentials were delegated to the remote peer.

• False - No credentials were delegated.

GSS_C_MUTUAL_FLAG

• True - The remote peer has authenticated itself.

• False - Remote peer has not authenticated itself.

GSS_C_REPLAY_FLAG

• True - replay of protected messages will be detected.

• False - replayed messages will not be detected.

GSS_C_SEQUENCE_FLAG

• True - out-of-sequence protected messages will be detected.

• False - out-of-sequence messages will not be detected.

GSS_C_CONF_FLAG

• True - Confidentiality service may be invoked by calling gss wrap routine.

• False - No confidentiality service (via gss wrap) available. gss wrap will provide message encapsulation, data-origin authentication and integrity services only.

GSS_C_INTEG_FLAG

• True - Integrity service may be invoked by calling either gss get mic or gss wrap routines.

• False - Per-message integrity service unavailable.

GSS_C_ANON_FLAG

• True - The initiator’s identity has not been revealed, and will not be revealed if any emitted token is passed to the acceptor.

• False - The initiator’s identity has been or will be authenticated normally.

          GSS_C_PROT_READY_FLAG

• True - Protection services (as specified by the states of the GSS C CONF FLAG and GSS C INTEG FLAG) are available for use if the accompanying major status return value is either GSS S COMPLETE or GSS S CONTINUE NEEDED.

• False - Protection services (as specified by the states of the GSS C CONF FLAG and GSS C INTEG FLAG) are available only if the accompanying major status return value is GSS S COMPLETE.

          GSS_C_TRANS_FLAG

• True - The resultant security context may be transferred to other processes via a call to gss export sec context().

• False - The security context is not transferable.

All other bits should be set to zero. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_CONTINUE_NEEDED: Indicates that a token from the peer application is required to complete the context, and that gss init sec context must be called again with that token. GSS_S_DEFECTIVE_TOKEN: Indicates that consistency checks performed on the input token failed. GSS_S_DEFECTIVE_CREDENTIAL: Indicates that consistency checks performed on the credential failed. GSS_S_NO_CRED: The supplied credentials were not valid for context initiation, or the credential handle did not reference any credentials. GSS_S_CREDENTIALS_EXPIRED: The referenced credentials have expired. GSS_S_BAD_BINDINGS: The input token contains different channel bindings to those specified via the input chan bindings parameter. GSS_S_BAD_SIG: The input token contains an invalid MIC, or a MIC that could not be verified. GSS_S_OLD_TOKEN: The input token was too old. This is a fatal error during context establishment. GSS_S_DUPLICATE_TOKEN: The input token is valid, but is a duplicate of a token already processed. This is a fatal error during context establishment. GSS_S_NO_CONTEXT: Indicates that the supplied context handle did not refer to a valid context. GSS_S_BAD_NAMETYPE: The provided target name parameter contained an invalid or unsupported type of name.

GSS_S_BAD_NAME: The provided target name parameter was ill-formed.

GSS_S_BAD_MECH: The specified mechanism is not supported by the provided credential, or is unrecognized by the implementation.

gss accept sec context

[Function] OM_uint32 gss_accept_sec_context (OM uint32 * minor_status,

gss ctx id t * context_handle, const gss cred id t

   acceptor_cred_handle, const gss buffer t input_token_buffer, const

gss channel bindings t input_chan_bindings, gss name t * src_name,

   gss OID * mech_type, gss buffer t output_token, OM uint32 * ret_flags,
   OM uint32 * time_rec, gss cred id t * delegated_cred_handle)

minor status: (Integer, modify) Mechanism specific status code.

context handle: (gss ctx id t, read/modify) Context handle for new context. Supply GSS C NO CONTEXT for first call; use value returned in subsequent calls. Once gss accept sec context() has returned a value via this parameter, resources have been assigned to the corresponding context, and must be freed by the application after use with a call to gss delete sec context().

acceptor cred handle: (gss cred id t, read) Credential handle claimed by context acceptor. Specify GSS C NO CREDENTIAL to accept the context as a default principal. If GSS C NO CREDENTIAL is specified, but no default acceptor principal is defined, GSS S NO CRED will be returned.

input token buffer: (buffer, opaque, read) Token obtained from remote application.

input chan bindings: (channel bindings, read, optional) Application- specified bindings. Allows application to securely bind channel identification information to the security context. If channel bindings are not used, specify GSS C NO CHANNEL BINDINGS.

src name: (gss name t, modify, optional) Authenticated name of context initiator. After use, this name should be deallocated by passing it to gss release name(). If not required, specify NULL.

mech type: (Object ID, modify, optional) Security mechanism used. The returned OID value will be a pointer into static storage, and should be treated as read-only by the caller (in particular, it does not need to be freed). If not required, specify NULL.

output token: (buffer, opaque, modify) Token to be passed to peer application. If the length field of the returned token buffer is 0, then no token need be passed to the peer application. If a non- zero length field is returned, the associated storage must be freed after use by the application with a call to gss release buffer().

ret flags: (bit-mask, modify, optional) Contains various independent flags, each of which indicates that the context supports a specific service option. If not needed, specify NULL. Symbolic names are provided for each flag, and the symbolic names corresponding to the required flags should be logically-ANDed with the ret flags value to test whether a given option is supported by the context. See below for the flags.

time rec: (Integer, modify, optional) Number of seconds for which the context will remain valid. Specify NULL if not required. delegated cred handle: (gss cred id t, modify, optional credential) Handle for credentials received from context initiator. Only valid if deleg flag in ret flags is true, in which case an explicit credential handle (i.e. not GSS C NO CREDENTIAL) will be returned; if deleg flag is false, gss accept sec context() will set this parameter to GSS C NO CREDENTIAL. If a credential handle is returned, the associated resources must be released by the application after use with a call to gss release cred(). Specify NULL if not required.

Allows a remotely initiated security context between the application and a remote peer to be established. The routine may return a output token which should be transferred to the peer application, where the peer application will present it to gss init sec context. If no token need be sent, gss accept sec context will indicate this by setting the length field of the output token argument to zero. To complete the context establishment, one or more reply tokens may be required from the peer application; if so, gss accept sec context will return a status flag of GSS S CONTINUE NEEDED, in which case it should be called again when the reply token is received from the peer application, passing the token to gss accept sec context via the input token parameters.

Portable applications should be constructed to use the token length and return status to determine whether a token needs to be sent or waited for. Thus a typical portable caller should always invoke gss accept sec context within a loop:

     gss_ctx_id_t context_hdl = GSS_C_NO_CONTEXT;
     do {
       receive_token_from_peer(input_token);
       maj_stat = gss_accept_sec_context(&min_stat,
                                         &context_hdl,
                                         cred_hdl,
                                         input_token,
                                         input_bindings,
                                         &client_name,
                                         &mech_type,
                                         output_token,
                                         &ret_flags,
                                         &time_rec,
                                         &deleg_cred);
       if (GSS_ERROR(maj_stat)) {
         report_error(maj_stat, min_stat);
       };
       if (output_token->length != 0) {
         send_token_to_peer(output_token);
         gss_release_buffer(&min_stat, output_token);
       };
       if (GSS_ERROR(maj_stat)) {
         if (context_hdl != GSS_C_NO_CONTEXT)
           gss_delete_sec_context(&min_stat,

                                  &context_hdl,
                                  GSS_C_NO_BUFFER);
         break;
       };
     } while (maj_stat & GSS_S_CONTINUE_NEEDED);

Whenever the routine returns a major status that includes the value GSS S CONTINUE NEEDED, the context is not fully established and the following restrictions apply to the output parameters: The value returned via the time rec parameter is undefined Unless the accompanying ret flags parameter contains the bit GSS C PROT READY FLAG, indicating that per-message services may be applied in advance of a successful completion status, the value returned via the mech type parameter may be undefined until the routine returns a major status value of GSS S COMPLETE. The values of the GSS C DELEG FLAG, GSS C MUTUAL FLAG,GSS C REPLAY FLAG, GSS C SEQUENCE FLAG, GSS C CONF FLAG,GSS C INTEG FLAG and GSS C ANON FLAG bits returned via the ret flags parameter should contain the values that the implementation expects would be valid if context establishment were to succeed. The values of the GSS C PROT READY FLAG and GSS C TRANS FLAG bits within ret flags should indicate the actual state at the time gss accept sec context returns, whether or not the context is fully established. Although this requires that GSS-API implementations set the GSS C PROT READY FLAG in the final ret flags returned to a caller (i.e. when accompanied by a GSS S COMPLETE status code), applications should not rely on this behavior as the flag was not defined in Version 1 of the GSS-API. Instead, applications should be prepared to use per-message services after a successful context establishment, according to the GSS C INTEG FLAG and GSS C CONF FLAG values. All other bits within the ret flags argument should be set to zero. While the routine returns GSS S CONTINUE NEEDED, the values returned via the ret flags argument indicate the services that the implementation expects to be available from the established context. If the initial call of gss accept sec context() fails, the implementation should not create a context object, and should leave the value of the context handle parameter set to GSS C NO CONTEXT to indicate this. In the event of a failure on a subsequent call, the implementation is permitted to delete the "half-built" security context (in which case it should set the context handle parameter to GSS C NO CONTEXT), but the preferred behavior is to leave the security context (and the context handle parameter) untouched for the application to delete (using gss delete sec context). During context establishment, the informational status bits GSS S OLD TOKEN and GSS S DUPLICATE TOKEN indicate fatal errors, and GSS-API mechanisms should always return them in association with a routine error of GSS S FAILURE. This requirement for pairing did not exist in version 1 of the GSS-API specification, so applications that wish to run over version 1 implementations must special-case these codes. The ret_flags values:

GSS_C_DELEG_FLAG

• True - Delegated credentials are available via the delegated cred handle parameter.

• False - No credentials were delegated.

GSS_C_MUTUAL_FLAG

• True - Remote peer asked for mutual authentication.

• False - Remote peer did not ask for mutual authentication.

GSS_C_REPLAY_FLAG

• True - replay of protected messages will be detected.

• False - replayed messages will not be detected.

GSS_C_SEQUENCE_FLAG

• True - out-of-sequence protected messages will be detected.

• False - out-of-sequence messages will not be detected.

GSS_C_CONF_FLAG

• True - Confidentiality service may be invoked by calling the gss wrap routine.

• False - No confidentiality service (via gss wrap) available. gss wrap will provide message encapsulation, data-origin authentication and integrity services only.

GSS_C_INTEG_FLAG

• True - Integrity service may be invoked by calling either gss get mic or gss wrap routines.

• False - Per-message integrity service unavailable.

GSS_C_ANON_FLAG

• True - The initiator does not wish to be authenticated; the src name parameter (if requested) contains an anonymous internal name.

• False - The initiator has been authenticated normally.

GSS_C_PROT_READY_FLAG

• True - Protection services (as specified by the states of the GSS C CONF FLAG and GSS C INTEG FLAG) are available if the accompanying major status return value is either GSS S COMPLETE or GSS S CONTINUE NEEDED.

• False - Protection services (as specified by the states of the GSS C CONF FLAG and GSS C INTEG FLAG) are available only if the accompanying major status return value is GSS S COMPLETE.

GSS_C_TRANS_FLAG

• True - The resultant security context may be transferred to other processes via a call to gss export sec context().

• False - The security context is not transferable.

All other bits should be set to zero. Return value: GSS_S_CONTINUE_NEEDED: Indicates that a token from the peer application is required to complete the context, and that gss accept sec context must be called again with that token. GSS_S_DEFECTIVE_TOKEN: Indicates that consistency checks performed on the input token failed. GSS_S_DEFECTIVE_CREDENTIAL: Indicates that consistency checks performed on the credential failed. GSS_S_NO_CRED: The supplied credentials were not valid for context acceptance, or the credential handle did not reference any credentials. GSS_S_CREDENTIALS_EXPIRED: The referenced credentials have expired. GSS_S_BAD_BINDINGS: The input token contains different channel bindings to those specified via the input chan bindings parameter. GSS_S_NO_CONTEXT: Indicates that the supplied context handle did not refer to a valid context. GSS_S_BAD_SIG: The input token contains an invalid MIC. GSS_S_OLD_TOKEN: The input token was too old. This is a fatal error during context establishment. GSS_S_DUPLICATE_TOKEN: The input token is valid, but is a duplicate of a token already processed. This is a fatal error during context establishment. GSS_S_BAD_MECH: The received token specified a mechanism that is not supported by the implementation or the provided credential.

gss delete sec context

[Function] OM_uint32 gss_delete_sec_context (OM uint32 * minor_status,
   gss ctx id t * context_handle, gss buffer t output_token)

minor status: (Integer, modify) Mechanism specific status code. context handle: (gss ctx id t, modify) Context handle identifying context to delete. After deleting the context, the GSS-API will set this context handle to GSS C NO CONTEXT. output token: (buffer, opaque, modify, optional) Token to be sent to remote application to instruct it to also delete the context. It is recommended that applications specify GSS C NO BUFFER for this parameter, requesting local deletion only. If a buffer parameter is provided by the application, the mechanism may return a token in it; mechanisms that implement only local deletion should set the length field of this token to zero to indicate to the application that no token is to be sent to the peer. Delete a security context. gss delete sec context will delete the local data structures associated with the specified security context, and may generate an output token, which when passed to the peer gss process context token will instruct it to do likewise. If no token is required by the mechanism, the GSS-API should set the length field of the output token (if provided) to zero. No further security services may be obtained using the context specified by context handle.

In addition to deleting established security contexts, gss delete sec context must also be able to delete "half-built" security contexts resulting from an incomplete sequence of gss init sec context()/gss accept sec context() calls.

The output token parameter is retained for compatibility with version 1 of the GSS-API. It is recommended that both peer applications invoke gss delete sec context passing the value GSS C NO BUFFER for the output token parameter, indicating that no token is required, and that gss delete sec context should simply delete local context data structures. If the application does pass a valid buffer to gss delete sec context, mechanisms are encouraged to return a zero-length token, indicating that no peer action is necessary, and that no token should be transferred by the application.

Return value:

GSS_S_COMPLETE: Successful completion.

GSS_S_NO_CONTEXT: No valid context was supplied.

gss process context token

[Function] OM_uint32 gss_process_context_token (OM uint32 *

minor_status, const gss ctx id t context_handle, const gss buffer t

   token_buffer)

minor status: (Integer, modify) Implementation specific status code.

context handle: (gss ctx id t, read) Context handle of context on which token is to be processed

token buffer: (buffer, opaque, read) Token to process.

Provides a way to pass an asynchronous token to the security service. Most contextlevel tokens are emitted and processed synchronously by gss init sec context and gss accept sec context, and the application is informed as to whether further tokens are expected by the GSS C CONTINUE NEEDED major status bit. Occasionally, a mechanism may need to emit a context-level token at a point when the peer entity is not expecting a token. For example, the initiator’s final call to gss init sec context may emit a token and return a status of GSS S COMPLETE, but the acceptor’s call to gss accept sec context may fail. The acceptor’s mechanism may wish to send a token containing an error indication to the initiator, but the initiator is not expecting a token at this point, believing that the context is fully established. Gss process context token provides a way to pass such a token to the mechanism at any time.

Return value:

GSS_S_COMPLETE: Successful completion.

GSS_S_DEFECTIVE_TOKEN: Indicates that consistency checks performed on the token failed.

GSS_S_NO_CONTEXT: The context handle did not refer to a valid context.

gss context time

[Function] OM_uint32 gss_context_time (OM uint32 * minor_status, const
   gss ctx id t context_handle, OM uint32 * time_rec)

minor status: (Integer, modify) Implementation specific status code. context handle: (gss ctx id t, read) Identifies the context to be interrogated. time rec: (Integer, modify) Number of seconds that the context will remain valid. If the context has already expired, zero will be returned. Determines the number of seconds for which the specified context will remain valid. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_CONTEXT_EXPIRED: The context has already expired. GSS_S_NO_CONTEXT: The context handle parameter did not identify a valid context

gss inquire context

[Function] OM_uint32 gss_inquire_context (OM uint32 * minor_status, const

gss ctx id t context_handle, gss name t * src_name, gss name t *

   targ_name, OM uint32 * lifetime_rec, gss OID * mech_type, OM uint32
   * ctx_flags, int * locally_initiated, int * open)

minor status: (Integer, modify) Mechanism specific status code. context handle: (gss ctx id t, read) A handle that refers to the security context. src name: (gss name t, modify, optional) The name of the context initiator. If the context was established using anonymous authentication, and if the application invoking gss inquire context is the context acceptor, an anonymous name will be returned. Storage associated with this name must be freed by the application after use with a call to gss release name(). Specify NULL if not required. targ name: (gss name t, modify, optional) The name of the context acceptor. Storage associated with this name must be freed by the application after use with a call to gss release name(). If the context acceptor did not authenticate itself, and if the initiator did not specify a target name in its call to gss init sec context(), the value GSS C NO NAME will be returned. Specify NULL if not required. lifetime rec: (Integer, modify, optional) The number of seconds for which the context will remain valid. If the context has expired, this parameter will be set to zero. If the implementation does not support context expiration, the value GSS C INDEFINITE will be returned. Specify NULL if not required. mech type: (gss OID, modify, optional) The security mechanism providing the context. The returned OID will be a pointer to static storage that should be treated as read-only by the application; in particular the application should not attempt to free it. Specify NULL if not required. ctx flags: (bit-mask, modify, optional) Contains various independent flags, each of which indicates that the context supports (or is expected to support, if ctx open is false) a specific service option. If not needed, specify NULL. Symbolic names are provided for each flag, and the symbolic names corresponding to the required flags should be logically-ANDed with the ret flags value to test whether a given option is supported by the context. See below for the flags. locally initiated: (Boolean, modify) Non-zero if the invoking application is the context initiator. Specify NULL if not required. open: (Boolean, modify) Non-zero if the context is fully established; Zero if a contextestablishment token is expected from the peer application. Specify NULL if not required. Obtains information about a security context. The caller must already have obtained a handle that refers to the context, although the context need not be fully established. The ctx_flags values:

GSS_C_DELEG_FLAG

• True - Credentials were delegated from the initiator to the acceptor.

• False - No credentials were delegated.

GSS_C_MUTUAL_FLAG

• True - The acceptor was authenticated to the initiator.

• False - The acceptor did not authenticate itself.

GSS_C_REPLAY_FLAG

• True - replay of protected messages will be detected.

• False - replayed messages will not be detected.

GSS_C_SEQUENCE_FLAG

• True - out-of-sequence protected messages will be detected.

• False - out-of-sequence messages will not be detected.

GSS_C_CONF_FLAG

• True - Confidentiality service may be invoked by calling gss wrap routine.

• False - No confidentiality service (via gss wrap) available. gss wrap will provide message encapsulation, data-origin authentication and integrity services only.

GSS_C_INTEG_FLAG

• True - Integrity service may be invoked by calling either gss get mic or gss wrap routines.

• False - Per-message integrity service unavailable.

GSS_C_ANON_FLAG

• True - The initiator’s identity will not be revealed to the acceptor. The src name parameter (if requested) contains an anonymous internal name.

• False - The initiator has been authenticated normally.

GSS_C_PROT_READY_FLAG

• True - Protection services (as specified by the states of the GSS C CONF FLAG and GSS C INTEG FLAG) are available for use.

• False - Protection services (as specified by the states of the GSS C CONF FLAG and GSS C INTEG FLAG) are available only if the context is fully established (i.e. if the open parameter is non-zero).

          GSS_C_TRANS_FLAG

• True - The resultant security context may be transferred to other processes via a call to gss export sec context().

• False - The security context is not transferable.

Return value: GSS_S_COMPLETE: Successful completion. GSS_S_NO_CONTEXT: The referenced context could not be accessed.

gss wrap size limit

[Function] OM_uint32 gss_wrap_size_limit (OM uint32 * minor_status, const
   gss ctx id t context_handle, int conf_req_flag, gss qop t qop_req,
   OM uint32 req_output_size, OM uint32 * max_input_size)

minor status: (Integer, modify) Mechanism specific status code. context handle: (gss ctx id t, read) A handle that refers to the security over which the messages will be sent. conf req flag: (Boolean, read) Indicates whether gss wrap will be asked to apply confidentiality protection in addition to integrity protection. See the routine description for gss wrap for more details. qop req: (gss qop t, read) Indicates the level of protection that gss wrap will be asked to provide. See the routine description for gss wrap for more details. req output size: (Integer, read) The desired maximum size for tokens emitted by gss wrap. max input size: (Integer, modify) The maximum input message size that may be presented to gss wrap in order to guarantee that the emitted token shall be no larger than req output size bytes. Allows an application to determine the maximum message size that, if presented to gss wrap with the same conf req flag and qop req parameters, will result in an output token containing no more than req output size bytes. This call is intended for use by applications that communicate over protocols that impose a maximum message size. It enables the application to fragment messages prior to applying protection. GSS-API implementations are recommended but not required to detect invalid QOP values when gss wrap size limit() is called. This routine guarantees only a maximum message size, not the availability of specific QOP values for message protection. Successful completion of this call does not guarantee that gss wrap will be able to protect a message of length max input size bytes, since this ability may depend on the availability of system resources at the time that gss wrap is called. However, if the implementation itself imposes an upper limit on the length of messages that may be processed by gss wrap, the implementation should not return a value via max input bytes that is greater than this length. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_NO_CONTEXT: The referenced context could not be accessed. GSS_S_CONTEXT_EXPIRED: The context has expired. GSS_S_BAD_QOP: The specified QOP is not supported by the mechanism.

gss export sec context

[Function] OM_uint32 gss_export_sec_context (OM uint32 * minor_status,
   gss ctx id t * context_handle, gss buffer t interprocess_token)

minor status: (Integer, modify) Mechanism specific status code. context handle: (gss ctx id t, modify) Context handle identifying the context to transfer. interprocess token: (buffer, opaque, modify) Token to be transferred to target process. Storage associated with this token must be freed by the application after use with a call to gss release buffer(). Provided to support the sharing of work between multiple processes. This routine will typically be used by the context-acceptor, in an application where a single process receives incoming connection requests and accepts security contexts over them, then passes the established context to one or more other processes for message exchange. gss export sec context() deactivates the security context for the calling process and creates an interprocess token which, when passed to gss import sec context in another process, will re-activate the context in the second process. Only a single instantiation of a given context may be active at any one time; a subsequent attempt by a context exporter to access the exported security context will fail. The implementation may constrain the set of processes by which the interprocess token may be imported, either as a function of local security policy, or as a result of implementation decisions. For example, some implementations may constrain contexts to be passed only between processes that run under the same account, or which are part of the same process group. The interprocess token may contain security-sensitive information (for example cryptographic keys). While mechanisms are encouraged to either avoid placing such sensitive information within interprocess tokens, or to encrypt the token before returning it to the application, in a typical object-library GSS-API implementation this may not be possible. Thus the application must take care to protect the interprocess token, and ensure that any process to which the token is transferred is trustworthy. If creation of the interprocess token is successful, the implementation shall deallocate all process-wide resources associated with the security context, and set the context handle to GSS C NO CONTEXT. In the event of an error that makes it impossible to complete the export of the security context, the implementation must not return an interprocess token, and should strive to leave the security context referenced by the context handle parameter untouched. If this is impossible, it is permissible for the implementation to delete the security context, providing it also sets the context handle parameter to GSS C NO CONTEXT. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_CONTEXT_EXPIRED: The context has expired. GSS_S_NO_CONTEXT: The context was invalid. GSS_S_UNAVAILABLE: The operation is not supported.

gss import sec context

[Function] OM_uint32 gss_import_sec_context (OM uint32 * minor_status,
   const gss buffer t interprocess_token, gss ctx id t * context_handle)

minor status: (Integer, modify) Mechanism specific status code. interprocess token: (buffer, opaque, modify) Token received from exporting process context handle: (gss ctx id t, modify) Context handle of newly reactivated context. Resources associated with this context handle must be released by the application after use with a call to gss delete sec context(). Allows a process to import a security context established by another process. A given interprocess token may be imported only once. See gss export sec context. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_NO_CONTEXT: The token did not contain a valid context reference.

GSS_S_DEFECTIVE_TOKEN: The token was invalid.

GSS_S_UNAVAILABLE: The operation is unavailable. GSS_S_UNAUTHORIZED: Local policy prevents the import of this context by the current process.

3.7 Per-Message Routines

GSS-API Per-message Routines
Routine                         Function
-------                         --------
gss_get_mic                     Calculate a cryptographic message
                              integrity code (MIC) for a
                              message; integrity service.
gss_verify_mic                  Check a MIC against a message;
                              verify integrity of a received
                              message.
gss_wrap                        Attach a MIC to a message, and
                              optionally encrypt the message
                              content.
                              confidentiality service
gss_unwrap                      Verify a message with attached

                              MIC, and decrypt message content
                              if necessary.

gss get mic

[Function] OM_uint32 gss_get_mic (OM uint32 * minor_status, const

gss ctx id t context_handle, gss qop t qop_req, const gss buffer t

   message_buffer, gss buffer t message_token)

minor status: (Integer, modify) Mechanism specific status code. context handle: (gss ctx id t, read) Identifies the context on which the message will be sent. qop req: (gss qop t, read, optional) Specifies requested quality of protection. Callers are encouraged, on portability grounds, to accept the default quality of protection offered by the chosen mechanism, which may be requested by specifying GSS C QOP DEFAULT for this parameter. If an unsupported protection strength is requested, gss get mic will return a major status of GSS S BAD QOP. message buffer: (buffer, opaque, read) Message to be protected. message token: (buffer, opaque, modify) Buffer to receive token. The application must free storage associated with this buffer after use with a call to gss release buffer(). Generates a cryptographic MIC for the supplied message, and places the MIC in a token for transfer to the peer application. The qop req parameter allows a choice between several cryptographic algorithms, if supported by the chosen mechanism. Since some application-level protocols may wish to use tokens emitted by gss wrap() to provide "secure framing", implementations must support derivation of MICs from zero-length messages. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_CONTEXT_EXPIRED: The context has already expired. GSS_S_NO_CONTEXT: The context handle parameter did not identify a valid context. GSS_S_BAD_QOP: The specified QOP is not supported by the mechanism.

gss verify mic

[Function] OM_uint32 gss_verify_mic (OM uint32 * minor_status, const

gss ctx id t context_handle, const gss buffer t message_buffer, const

   gss buffer t token_buffer, gss qop t * qop_state)

minor status: (Integer, modify) Mechanism specific status code. context handle: (gss ctx id t, read) Identifies the context on which the message arrived. message buffer: (buffer, opaque, read) Message to be verified. token buffer: (buffer, opaque, read) Token associated with message. qop state: (gss qop t, modify, optional) Quality of protection gained from MIC Specify NULL if not required.

Verifies that a cryptographic MIC, contained in the token parameter, fits the supplied message. The qop state parameter allows a message recipient to determine the strength of protection that was applied to the message. Since some application-level protocols may wish to use tokens emitted by gss wrap() to provide "secure framing", implementations must support the calculation and verification of MICs over zero-length messages. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_DEFECTIVE_TOKEN: The token failed consistency checks. GSS_S_BAD_SIG: The MIC was incorrect. GSS_S_DUPLICATE_TOKEN: The token was valid, and contained a correct MIC for the message, but it had already been processed. GSS_S_OLD_TOKEN: The token was valid, and contained a correct MIC for the message, but it is too old to check for duplication. GSS_S_UNSEQ_TOKEN: The token was valid, and contained a correct MIC for the message, but has been verified out of sequence; a later token has already been received. GSS_S_GAP_TOKEN: The token was valid, and contained a correct MIC for the message, but has been verified out of sequence; an earlier expected token has not yet been received. GSS_S_CONTEXT_EXPIRED: The context has already expired. GSS_S_NO_CONTEXT: The context handle parameter did not identify a valid context.

gss wrap

[Function] OM_uint32 gss_wrap (OM uint32 * minor_status, const gss ctx id t

   context_handle, int conf_req_flag, gss qop t qop_req, const gss buffer t
   input_message_buffer, int * conf_state, gss buffer t
   output_message_buffer)

minor status: (Integer, modify) Mechanism specific status code. context handle: (gss ctx id t, read) Identifies the context on which the message will be sent. conf req flag: (boolean, read) Non-zero - Both confidentiality and integrity services are requested. Zero - Only integrity service is requested. qop req: (gss qop t, read, optional) Specifies required quality of protection. A mechanism-specific default may be requested by setting qop req to GSS C QOP DEFAULT. If an unsupported protection strength is requested, gss wrap will return a major status of GSS S BAD QOP. input message buffer: (buffer, opaque, read) Message to be protected. conf state: (boolean, modify, optional) Non-zero - Confidentiality, data origin authentication and integrity services have been applied. Zero - Integrity and data origin services only has been applied. Specify NULL if not required. output message buffer: (buffer, opaque, modify) Buffer to receive protected message. Storage associated with this message must be freed by the application after use with a call to gss release buffer().

Attaches a cryptographic MIC and optionally encrypts the specified input message. The output message contains both the MIC and the message. The qop req parameter allows a choice between several cryptographic algorithms, if supported by the chosen mechanism. Since some application-level protocols may wish to use tokens emitted by gss wrap() to provide "secure framing", implementations must support the wrapping of zerolength messages. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_CONTEXT_EXPIRED: The context has already expired. GSS_S_NO_CONTEXT: The context handle parameter did not identify a valid context. GSS_S_BAD_QOP: The specified QOP is not supported by the mechanism.

gss unwrap

[Function] OM_uint32 gss_unwrap (OM uint32 * minor_status, const
   gss ctx id t context_handle, const gss buffer t input_message_buffer,
   gss buffer t output_message_buffer, int * conf_state, gss qop t *
   qop_state)

minor status: (Integer, modify) Mechanism specific status code. context handle: (gss ctx id t, read) Identifies the context on which the message arrived. input message buffer: (buffer, opaque, read) Protected message. output message buffer: (buffer, opaque, modify) Buffer to receive unwrapped message. Storage associated with this buffer must be freed by the application after use use with a call to gss release buffer(). conf state: (boolean, modify, optional) Non-zero - Confidentiality and integrity protection were used. Zero - Integrity service only was used. Specify NULL if not required. qop state: (gss qop t, modify, optional) Quality of protection provided. Specify NULL if not required. Converts a message previously protected by gss wrap back to a usable form, verifying the embedded MIC. The conf state parameter indicates whether the message was encrypted; the qop state parameter indicates the strength of protection that was used to provide the confidentiality and integrity services. Since some application-level protocols may wish to use tokens emitted by gss wrap() to provide "secure framing", implementations must support the wrapping and unwrapping of zero-length messages. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_DEFECTIVE_TOKEN: The token failed consistency checks. GSS_S_BAD_SIG: The MIC was incorrect. GSS_S_DUPLICATE_TOKEN: The token was valid, and contained a correct MIC for the message, but it had already been processed.

GSS_S_OLD_TOKEN: The token was valid, and contained a correct MIC for the message, but it is too old to check for duplication.

GSS_S_UNSEQ_TOKEN: The token was valid, and contained a correct MIC for the message, but has been verified out of sequence; a later token has already been received.

GSS_S_GAP_TOKEN: The token was valid, and contained a correct MIC for the message, but has been verified out of sequence; an earlier expected token has not yet been received.

GSS_S_CONTEXT_EXPIRED: The context has already expired.

GSS_S_NO_CONTEXT: The context handle parameter did not identify a valid context.

3.8 Name Manipulation

GSS-API Name manipulation Routines
Routine                         Function
-------                         --------
gss_import_name                 Convert a contiguous string name
                              to internal-form.
gss_display_name                Convert internal-form name to
                              text.
gss_compare_name                Compare two internal-form names.
gss_release_name                Discard an internal-form name.
gss_inquire_names_for_mech      List the name-types supported by.
                              the specified mechanism.
gss_inquire_mechs_for_name      List mechanisms that support the
                              specified name-type.
gss_canonicalize_name           Convert an internal name to an MN.
gss_export_name                 Convert an MN to export form.
gss_duplicate_name              Create a copy of an internal name.

gss import name

[Function] OM_uint32 gss_import_name (OM uint32 * minor_status, const
   gss buffer t input_name_buffer, const gss OID input_name_type,
   gss name t * output_name)

minor status: (Integer, modify) Mechanism specific status code.

input name buffer: (buffer, octet-string, read) Buffer containing contiguous string name to convert.

input name type: (Object ID, read, optional) Object ID specifying type of printable name. Applications may specify either GSS C NO OID to use a mechanism-specific default printable syntax, or an OID recognized by the GSS-API implementation to name a specific namespace.

output name: (gss name t, modify) Returned name in internal form. Storage associated with this name must be freed by the application after use with a call to gss release name().

Convert a contiguous string name to internal form. In general, the internal name returned (via the @output name parameter) will not be an MN; the exception to this is if the @input name type indicates that the contiguous string provided via the @input name buffer parameter is of type GSS C NT EXPORT NAME, in which case the returned internal name will be an MN for the mechanism that exported the name.

Return value:

GSS_S_COMPLETE: Successful completion.

GSS_S_BAD_NAMETYPE: The input name type was unrecognized.

GSS_S_BAD_NAME: The input name parameter could not be interpreted as a name of the specified type.

GSS_S_BAD_MECH: The input name-type was GSS C NT EXPORT NAME, but the mechanism contained within the input-name is not supported.

gss display name

[Function] OM_uint32 gss_display_name (OM uint32 * minor_status, const
   gss name t input_name, gss buffer t output_name_buffer, gss OID *
   output_name_type)

minor status: (Integer, modify) Mechanism specific status code.

input name: (gss name t, read) Name to be displayed.

output name buffer: (buffer, character-string, modify) Buffer to receive textual name string. The application must free storage associated with this name after use with a call to gss release buffer().

output name type: (Object ID, modify, optional) The type of the returned name. The returned gss OID will be a pointer into static storage, and should be treated as read-only by the caller (in particular, the application should not attempt to free it). Specify NULL if not required.

Allows an application to obtain a textual representation of an opaque internal-form name for display purposes. The syntax of a printable name is defined by the GSS-API implementation.

If input name denotes an anonymous principal, the implementation should return the gss OID value GSS C NT ANONYMOUS as the output name type, and a textual name that is syntactically distinct from all valid supported printable names in output name buffer.

If input name was created by a call to gss import name, specifying GSS C NO OID as the name-type, implementations that employ lazy conversion between name types may return GSS C NO OID via the output name type parameter.

Return value:

GSS_S_COMPLETE: Successful completion.

GSS_S_BAD_NAME: @input name was ill-formed.

gss compare name

[Function] OM_uint32 gss_compare_name (OM uint32 * minor_status, const

gss name t name1, const gss name t name2, int * name_equal) minor status: (Integer, modify) Mechanism specific status code. name1: (gss name t, read) Internal-form name. name2: (gss name t, read) Internal-form name. name equal: (boolean, modify) Non-zero - names refer to same entity. Zero - names refer to different entities (strictly, the names are not known to refer to the same identity). Allows an application to compare two internal-form names to determine whether they refer to the same entity. If either name presented to gss compare name denotes an anonymous principal, the routines should indicate that the two names do not refer to the same identity. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_BAD_NAMETYPE: The two names were of incomparable types. GSS_S_BAD_NAME: One or both of name1 or name2 was ill-formed.

gss release name

[Function] OM_uint32 gss_release_name (OM uint32 * minor_status,

gss name t * name) minor status: (Integer, modify) Mechanism specific status code. name: (gss name t, modify) The name to be deleted. Free GSSAPI-allocated storage associated with an internal-form name. The name is set to GSS C NO NAME on successful completion of this call. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_BAD_NAME: The name parameter did not contain a valid name.

gss inquire names for mech

[Function] OM_uint32 gss_inquire_names_for_mech (OM uint32 *
   minor_status, const gss OID mechanism, gss OID set * name_types)

minor status: (Integer, modify) Mechanism specific status code. mechanism: (gss OID, read) The mechanism to be interrogated. name types: (gss OID set, modify) Set of name-types supported by the specified mechanism. The returned OID set must be freed by the application after use with a call to gss release oid set(). Returns the set of nametypes supported by the specified mechanism. Return value: GSS_S_COMPLETE: Successful completion.

gss inquire mechs for name

[Function] OM_uint32 gss_inquire_mechs_for_name (OM uint32 *
   minor_status, const gss name t input_name, gss OID set * mech_types)

minor status: (Integer, modify) Mechanism specific status code. input name: (gss name t, read) The name to which the inquiry relates. mech types: (gss OID set, modify) Set of mechanisms that may support the specified name. The returned OID set must be freed by the caller after use with a call to gss release oid set(). Returns the set of mechanisms supported by the GSS-API implementation that may be able to process the specified name. Each mechanism returned will recognize at least one element within the name. It is permissible for this routine to be implemented within a mechanism-independent GSS-API layer, using the type information contained within the presented name, and based on registration information provided by individual mechanism implementations. This means that the returned mech types set may indicate that a particular mechanism will understand the name when in fact it would refuse to accept the name as input to gss canonicalize name, gss init sec context, gss acquire cred or gss add cred (due to some property of the specific name, as opposed to the name type). Thus this routine should be used only as a prefilter for a call to a subsequent mechanism-specific routine. Return value: GSS_S_COMPLETE: Successful completion. GSS_S_BAD_NAME: The input name parameter was ill-formed. GSS_S_BAD_NAMETYPE: The input name parameter contained an invalid or unsupported type of name.

gss canonicalize name

[Function] OM_uint32 gss_canonicalize_name (OM uint32 * minor_status,

const gss name t input_name, const gss OID mech_type, gss name t *

   output_name)

minor status: (Integer, modify) Mechanism specific status code. input name: (gss name t, read) The name for which a canonical form is desired. mech type: (Object ID, read) The authentication mechanism for which the canonical form of the name is desired. The desired mechanism must be specified explicitly; no default is provided. output name: (gss name t, modify) The resultant canonical name. Storage associated with this name must be freed by the application after use with a call to gss release name(). Generate a canonical mechanism name (MN) from an arbitrary internal name. The mechanism name is the name that would be returned to a context acceptor on successful authentication of a context where the initiator used the input name in a successful call to gss acquire cred, specifying an OID set containing @mech type as its only member, followed by a call to gss init sec context(), specifying @mech type as the authentication mechanism.

Return value: GSS_S_COMPLETE: Successful completion.

gss export name

[Function] OM_uint32 gss_export_name (OM uint32 * minor_status, const
   gss name t input_name, gss buffer t exported_name)

minor status: (Integer, modify) Mechanism specific status code. input name: (gss name t, read) The MN to be exported. exported name: (gss buffer t, octet-string, modify) The canonical contiguous string form of @input name. Storage associated with this string must freed by the application after use with gss release buffer(). To produce a canonical contiguous string representation of a mechanism name (MN), suitable for direct comparison (e.g. with memcmp) for use in authorization functions (e.g. matching entries in an access-control list). The @input name parameter must specify a valid MN (i.e. an internal name generated by gss accept sec context() or by gss canonicalize name()). Return value: GSS_S_COMPLETE: Successful completion. GSS_S_NAME_NOT_MN: The provided internal name was not a mechanism name. GSS_S_BAD_NAME: The provided internal name was ill-formed. GSS_S_BAD_NAMETYPE: The internal name was of a type not supported by the GSS-API implementation.

gss duplicate name

[Function] OM_uint32 gss_duplicate_name (OM uint32 * minor_status, const

gss name t src_name, gss name t * dest_name) minor status: (Integer, modify) Mechanism specific status code. src name: (gss name t, read) Internal name to be duplicated. dest name: (gss name t, modify) The resultant copy of @src name. Storage associated with this name must be freed by the application after use with a call to gss release name(). Create an exact duplicate of the existing internal name @src name. The new @dest name will be independent of src name (i.e. @src name and @dest name must both be released, and the release of one shall not affect the validity of the other). Return value: GSS_S_COMPLETE: Successful completion. GSS_S_BAD_NAME: The src name parameter was ill-formed.

3.9 Miscellaneous Routines

GSS-API Miscellaneous Routines
Routine                        Function
-------                        --------
gss_add_oid_set_member         Add an object identifier to
                             a set.
gss_display_status             Convert a GSS-API status code
                             to text.
gss_indicate_mechs             Determine available underlying
                             authentication mechanisms.
gss_release_buffer             Discard a buffer.
gss_release_oid_set            Discard a set of object
                             identifiers.
gss_create_empty_oid_set       Create a set containing no
                             object identifiers.
gss_test_oid_set_member        Determines whether an object
                             identifier is a member of a set.
gss_encapsulate_token          Encapsulate a context token.
gss_decapsulate_token          Decapsulate a context token.
gss_oid_equal                  Compare two OIDs for equality.

gss add oid set member

[Function] OM_uint32 gss_add_oid_set_member (OM uint32 * minor_status,

const gss OID member_oid, gss OID set * oid_set) minor status: (integer, modify) Mechanism specific status code.

member oid: (Object ID, read) The object identifier to copied into the set.

oid set: (Set of Object ID, modify) The set in which the object identifier should be inserted.

Add an Object Identifier to an Object Identifier set. This routine is intended for use in conjunction with gss create empty oid set when constructing a set of mechanism OIDs for input to gss acquire cred. The oid set parameter must refer to an OIDset that was created by GSS-API (e.g. a set returned by gss create empty oid set()). GSS-API creates a copy of the member oid and inserts this copy into the set, expanding the storage allocated to the OID-set’s elements array if necessary. The routine may add the new member OID anywhere within the elements array, and implementations should verify that the new member oid is not already contained within the elements array; if the member oid is already present, the oid set should remain unchanged.

Return value:

GSS_S_COMPLETE: Successful completion.

gss display status

[Function] OM_uint32 gss_display_status (OM uint32 * minor_status,
   OM uint32 status_value, int status_type, const gss OID mech_type,
   OM uint32 * message_context, gss buffer t status_string)

minor status: (integer, modify) Mechanism specific status code. status value: (Integer, read) Status value to be converted. status type: (Integer, read) GSS C GSS CODE - status value is a GSS status code. GSS C MECH CODE - status value is a mechanism status code. mech type: (Object ID, read, optional) Underlying mechanism (used to interpret a minor status value). Supply GSS C NO OID to obtain the system default. message context: (Integer, read/modify) Should be initialized to zero by the application prior to the first call. On return from gss display status(), a non-zero status value parameter indicates that additional messages may be extracted from the status code via subsequent calls to gss display status(), passing the same status value, status type, mech type, and message context parameters. status string: (buffer, character string, modify) Textual interpretation of the status value. Storage associated with this parameter must be freed by the application after use with a call to gss release buffer(). Allows an application to obtain a textual representation of a GSS-API status code, for display to the user or for logging purposes. Since some status values may indicate multiple conditions, applications may need to call gss display status multiple times, each call generating a single text string. The message context parameter is used by gss display status to store state information about which error messages have already been extracted from a given status value; message context must be initialized to 0 by the application prior to the first call, and gss display status will return a non-zero value in this parameter if there are further messages to extract. The message context parameter contains all state information required by gss display status in order to extract further messages from the status value; even when a non-zero value is returned in this parameter, the application is not required to call gss display status again unless subsequent messages are desired. The following code extracts all messages from a given status code and prints them to stderr:

     OM_uint32 message_context;
     OM_uint32 status_code;
     OM_uint32 maj_status;
     OM_uint32 min_status;
     gss_buffer_desc status_string;
            ...
     message_context = 0;
     do {
       maj_status = gss_display_status (

                       &min_status,
                       status_code,
                       GSS_C_GSS_CODE,
                       GSS_C_NO_OID,
                       &message_context,
                       &status_string)
       fprintf(stderr,
               "%.*s\n",
              (int)status_string.length,
              (char *)status_string.value);
       gss_release_buffer(&min_status, &status_string);
     } while (message_context != 0);

Return value: GSS_S_COMPLETE: Successful completion. GSS_S_BAD_MECH: Indicates that translation in accordance with an unsupported mechanism type was requested. GSS_S_BAD_STATUS: The status value was not recognized, or the status type was neither GSS C GSS CODE nor GSS C MECH CODE.

gss indicate mechs

[Function] OM_uint32 gss_indicate_mechs (OM uint32 * minor_status,

gss OID set * mech_set) minor status: (integer, modify) Mechanism specific status code. mech set: (set of Object IDs, modify) Set of implementation-supported mechanisms. The returned gss OID set value will be a dynamically-allocated OID set, that should be released by the caller after use with a call to gss release oid set(). Allows an application to determine which underlying security mechanisms are available. Return value: GSS_S_COMPLETE: Successful completion.

gss release buffer

[Function] OM_uint32 gss_release_buffer (OM uint32 * minor_status,

gss buffer t buffer) minor status: (integer, modify) Mechanism specific status code. buffer: (buffer, modify) The storage associated with the buffer will be deleted. The gss buffer desc object will not be freed, but its length field will be zeroed. Free storage associated with a buffer. The storage must have been allocated by a GSS-API routine. In addition to freeing the associated storage, the routine will zero the length field in the descriptor to which the buffer parameter refers, and implementations are encouraged to additionally set the pointer field in the descriptor to NULL. Any buffer object returned by a GSS-API routine may be passed to gss release buffer (even if there is no storage associated with the buffer). Return value: GSS_S_COMPLETE: Successful completion.

gss release oid set

[Function] OM_uint32 gss_release_oid_set (OM uint32 * minor_status,

gss OID set * set) minor status: (integer, modify) Mechanism specific status code. set: (Set of Object IDs, modify) The storage associated with the gss OID set will be deleted. Free storage associated with a GSSAPI-generated gss OID set object. The set parameter must refer to an OID-set that was returned from a GSS-API routine. gss release oid set() will free the storage associated with each individual member OID, the OID set’s elements array, and the gss OID set desc. The gss OID set parameter is set to GSS C NO OID SET on successful completion of this routine. Return value: GSS_S_COMPLETE: Successful completion.

gss create empty oid set

[Function] OM_uint32 gss_create_empty_oid_set (OM uint32 *
   minor_status, gss OID set * oid_set)

minor status: (integer, modify) Mechanism specific status code. oid set: (Set of Object IDs, modify) The empty object identifier set. The routine will allocate the gss OID set desc object, which the application must free after use with a call to gss release oid set(). Create an object-identifier set containing no object identifiers, to which members may be subsequently added using the gss add oid set member() routine. These routines are intended to be used to construct sets of mechanism object identifiers, for input to gss acquire cred. Return value: GSS_S_COMPLETE: Successful completion.

gss test oid set member

[Function] OM_uint32 gss_test_oid_set_member (OM uint32 * minor_status,

const gss OID member, const gss OID set set, int * present) minor status: (integer, modify) Mechanism specific status code. member: (Object ID, read) The object identifier whose presence is to be tested. set: (Set of Object ID, read) The Object Identifier set. present: (Boolean, modify) Non-zero if the specified OID is a member of the set, zero if not. Interrogate an Object Identifier set to determine whether a specified Object Identifier is a member. This routine is intended to be used with OID sets returned by gss indicate mechs(), gss acquire cred(), and gss inquire cred(), but will also work with user-generated sets. Return value: GSS_S_COMPLETE: Successful completion.

gss encapsulate token

[Function] extern OM_uint32 gss_encapsulate_token (gss const buffer t
   input_token, gss const OID token_oid, gss buffer t output_token)

input token: (buffer, opaque, read) Buffer with GSS-API context token data. token oid: (Object ID, read) Object identifier of token. output token: (buffer, opaque, modify) Encapsulated token data; caller must release with gss release buffer(). Add the mechanism-independent token header to GSS-API context token data. This is used for the initial token of a GSS-API context establishment sequence. It incorporates an identifier of the mechanism type to be used on that context, and enables tokens to be interpreted unambiguously at GSS-API peers. See further section 3.1 of RFC 2743. This function is standardized in RFC 6339. Returns: GSS_S_COMPLETE: Indicates successful completion, and that output parameters holds correct information. GSS_S_FAILURE: Indicates that encapsulation failed for reasons unspecified at the GSS-API level.

gss decapsulate token

[Function] OM_uint32 gss_decapsulate_token (gss const buffer t input_token,

gss const OID token_oid, gss buffer t output_token) input token: (buffer, opaque, read) Buffer with GSS-API context token. token oid: (Object ID, read) Expected object identifier of token. output token: (buffer, opaque, modify) Decapsulated token data; caller must release with gss release buffer(). Remove the mechanism-independent token header from an initial GSS-API context token. Unwrap a buffer in the mechanism-independent token format. This is the reverse of gss encapsulate token(). The translation is loss-less, all data is preserved as is. This function is standardized in RFC 6339. Return value: GSS_S_COMPLETE: Indicates successful completion, and that output parameters holds correct information. GSS_S_DEFECTIVE_TOKEN: Means that the token failed consistency checks (e.g., OID mismatch or ASN.1 DER length errors).

GSS_S_FAILURE: Indicates that decapsulation failed for reasons unspecified at the GSS-API level.

gss oid equal

[Function] int gss_oid_equal (gss const OID first_oid, gss const OID

   second_oid)

first oid: (Object ID, read) First Object identifier. second oid: (Object ID, read) First Object identifier. Compare two OIDs for equality. The comparison is "deep", i.e., the actual byte sequences of the OIDs are compared instead of just the pointer equality. This function is standardized in RFC 6339. Return value: Returns boolean value true when the two OIDs are equal, otherwise false.

3.10 SASL GS2 Routines

gss inquire mech for saslname

[Function] OM_uint32 gss_inquire_mech_for_saslname (OM uint32 *
   minor_status, const gss buffer t sasl_mech_name, gss OID * mech_type)

minor status: (Integer, modify) Mechanism specific status code. sasl mech name: (buffer, character-string, read) Buffer with SASL mechanism name. mech type: (OID, modify, optional) Actual mechanism used. The OID returned via this parameter will be a pointer to static storage that should be treated as read-only; In particular the application should not attempt to free it. Specify NULL if not required. Output GSS-API mechanism OID of mechanism associated with given @sasl mech name. Returns: GSS_S_COMPLETE: Successful completion. GSS_S_BAD_MECH: There is no GSS-API mechanism known as @sasl mech name.

gss inquire saslname for mech

[Function] OM_uint32 gss_inquire_saslname_for_mech (OM uint32 *
   minor_status, const gss OID desired_mech, gss buffer t
   sasl_mech_name, gss buffer t mech_name, gss buffer t mech_description)

minor status: (Integer, modify) Mechanism specific status code. desired mech: (OID, read) Identifies the GSS-API mechanism to query. sasl mech name: (buffer, character-string, modify, optional) Buffer to receive SASL mechanism name. The application must free storage associated with this name after use with a call to gss release buffer(). mech name: (buffer, character-string, modify, optional) Buffer to receive human readable mechanism name. The application must free storage associated with this name after use with a call to gss release buffer(). mech description: (buffer, character-string, modify, optional) Buffer to receive description of mechanism. The application must free storage associated with this name after use with a call to gss release buffer(). Output the SASL mechanism name of a GSS-API mechanism. It also returns a name and description of the mechanism in a user friendly form. Returns: GSS_S_COMPLETE: Successful completion. GSS_S_BAD_MECH: The @desired mech OID is unsupported.

4 Extended GSS API

None of the following functions are standard GSS API functions. As such, they are not declared in gss/api.h, but rather in gss/ext.h (which is included from gss.h). See Section 2.1 [Header], page 7.

gss check version

[Function] const char * gss_check_version (const char * req_version)

req version: version string to compare with, or NULL Check that the version of the library is at minimum the one given as a string in @req version. Return value: The actual version string of the library; NULL if the condition is not met. If NULL is passed to this function no check is done and only the version string is returned.

gss userok

[Function] int gss_userok (const gss name t name, const char * username)

name: (gss name t, read) Name to be compared. username: Zero terminated string with username. Compare the username against the output from gss export name() invoked on @name, after removing the leading OID. This answers the question whether the particular mechanism would authenticate them as the same principal Return value: Returns 0 if the names match, non-0 otherwise.

5 Invoking gss

Name

GNU GSS (gss) – Command line interface to the GSS Library.

Description

gss is the main program of GNU GSS.

Mandatory or optional arguments to long options are also mandatory or optional for any corresponding short options.

Commands

gss recognizes these commands:

-l, --list-mechanisms
                  List information about supported mechanisms
                  in a human readable format.
-m, --major=LONG  Describe a `major status' error code value.
-a, --accept-sec-context
                  Accept a security context as server.
-i, --init-sec-context=MECH
                  Initialize a security context as client.
                  MECH is the SASL name of mechanism, use -l
                  to list supported mechanisms.
-n, --server-name=SERVICE@HOSTNAME
                  For -i, set the name of the remote host.
                  For example, "imap@mail.example.com".

Other Options

These are some standard parameters.

-h, --help        Print help and exit
-V, --version     Print version and exit
-q, --quiet       Silent operation  (default=off)

Examples

To list the supported mechanisms, use gss -l like this:

$ src/gss -l
Found 1 supported mechanisms.
Mechanism 0:
        Mechanism name: Kerberos V5
        Mechanism description: Kerberos V5 GSS-API mechanism
        SASL Mechanism name: GS2-KRB5
$

To initialize a Kerberos V5 security context, use the --init-sec-context parameter. Kerberos V5 needs to know the name of the remote entity, so you need to supply the --server-name parameter as well. That will provide the name of the server. For example, use imap@mail.example.com to setup a security context with the imap service on the host mail.example.com. The Kerberos V5 client will use your ticket-granting ticket (which needs to be available) and acquire a server ticket for the service. The KDC must know about the server for this to work. The tool will print the GSS-API context tokens base64 encoded on standard output.

$ gss -i GS2-KRB5 -n host@interop.josefsson.org
Context token (protection is available):
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
Input context token:

The tool is waiting for the final Kerberos V5 context token from the server. Note the status text informing you that message protection is available.

To accept a Kerberos V5 context, the process is similar. The server needs to know its name, so that it can find the host key from (typically) /etc/shishi/shishi.keys. Once started it will wait for a context token from the client. Below we’ll paste in the token printed above.

$ gss -a -n host@interop.josefsson.org
Importing name "host@interop.josefsson.org"...
Acquiring credentials...
Input context token:
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
Context has been accepted.  Final context token:
YHEGCSqGSIb3EgECAgIAb2IwYKADAgEFoQMCAQ+iVDBSoAMCARKhAwIBAKJGBESy1Zoy9DrG+DuV/6aWmAp79s9d+ofGXC/WKOzRuxAqo98vMRWbsbILW8z9aF1th4GZz0kjFz/hZAmnWyomZ9JiP3yQvg==
$

Returning to the client, you may now cut’n’paste the final context token as shown by the server. The client has then authenticated the server as well. The output from the client is shown below.

YHEGCSqGSIb3EgECAgIAb2IwYKADAgEFoQMCAQ+iVDBSoAMCARKhAwIBAKJGBESy1Zoy9DrG+DuV/6aWmAp79s9d+ofGXC/WKOzRuxAqo98vMRWbsbILW8z9aF1th4GZz0kjFz/hZAmnWyomZ9JiP3yQvg==
Context has been initialized.
$

Chapter 6: Acknowledgements 60

6 Acknowledgements

This manual borrows text from RFC 2743 and RFC 2744 that describe GSS API formally.

Appendix A Criticism of GSS

The author has doubts whether GSS is the best solution for free software projects looking for a implementation agnostic security framework. We express these doubts in this section, so that the reader can judge for herself if any of the potential problems discussed here are relevant for their project, or if the benefit outweigh the problems. We are aware that some of the opinions are highly subjective, but we offer them in the hope they can serve as anecdotal evidence.

GSS can be criticized on several levels. We start with the actual implementation.

GSS does not appear to be designed by experienced C programmers. While generally this may be a good thing (C is not the best language), but since they defined the API in C, it is unfortunate. The primary evidence of this is the major status and minor status error code solution. It is a complicated way to describe error conditions, but what makes matters worse, the error condition is separated; half of the error condition is in the function return value and the other half is in the first argument to the function, which is always a pointer to an integer. (The pointer is not even allowed to be NULL, if the application doesn’t care about the minor error code.) This makes the API unreadable, and difficult to use. A better solutions would be to return a struct containing the entire error condition, which can be accessed using macros, although we acknowledge that the C language used at the time GSS was designed may not have allowed this (this may in fact be the reason the awkward solution was chosen). Instead, the return value could have been passed back to callers using a pointer to a struct, accessible using various macros, and the function could have a void prototype. The fact that minor status is placed first in the parameter list increases the pain it is to use the API. Important parameters should be placed first. A better place for minor status (if it must be present at all) would have been last in the prototypes.

Another evidence of the C inexperience are the memory management issues; GSS provides functions to deallocate data stored within, e.g., gss_buffer_t but the caller is responsible of deallocating the structure pointed at by the gss_buffer_t (i.e., the gss_buffer_ desc) itself. Memory management issues are error prone, and this division easily leads to memory leaks (or worse). Instead, the API should be the sole owner of all gss_ctx_id_t, gss_cred_id_t, and gss_buffer_t structures: they should be allocated by the library, and deallocated (using the utility functions defined for this purpose) by the library.

TBA: specification is unclear how memory for OIDs are managed. For example, who is responsible for deallocate potentially newly allocated OIDs returned as actual_mechs in gss_acquire_cred? Further, are OIDs deeply copied into OID sets? In other words, if I add an OID into an OID set, and modify the original OID, will the OID in the OID set be modified too?

Another illustrating example is the sample GSS header file given in the RFC, which contains:

     /*
      * We have included the xom.h header file.  Verify that OM_uint32
      * is defined correctly.
      */
     #if sizeof(gss_uint32) != sizeof(OM_uint32)
     #error Incompatible definition of OM_uint32 from xom.h

     #endif

The C pre-processor does not know about the sizeof function, so it is treated as an identifier, which maps to 0. Thus, the expression does not check that the size of OM_uint32 is correct. It checks whether the expression 0 != 0 holds.

TBA: thread issues

TBA: multiple mechanisms in a GSS library

TBA: high-level design criticism.

TBA: no credential forwarding.

TBA: internationalization

TBA: dynamically generated OIDs and memory deallocation issue. I.e., should gss import name or gss duplicate name allocate memory and copy the OID provided, or simply copy the pointer? If the former, who would deallocate that memory? If the latter, the application may deallocate or modify the OID, which seem unwanted.

TBA: krb5: no way to access authorization-data

TBA: krb5: firewall/pre-IP: iakerb status?

TBA: krb5: single-DES only

TBA: the API may block, unusable in select() based servers. Especially if the servers contacted is decided by the, yet unauthenticated, remote client.

TBA: krb5: no support for GSS C PROT READY FLAG. We support it anyway, though.

TBA: krb5: gssapi-cfx differ from rfc 1964 in the reply token in that the latter require presence of sequence numbers whereas the former doesn’t.

Finally we note that few free security applications uses GSS, perhaps the only major exception to this are Kerberos 5 implementations. While not substantial evidence, this do suggest that the GSS may not be the simplest solution available to solve actual problems, since otherwise more projects would have chosen to take advantage of the work that went into GSS instead of using another framework (or designing their own solution).

Our conclusion is that free software projects that are looking for a security framework should evaluate carefully whether GSS actually is the best solution before using it. In particular it is recommended to compare GSS with the Simple Authentication and Security Layer (SASL) framework, which in several situations provide the same feature as GSS does. The most compelling argument for SASL over GSS is, as its acronym suggest, Simple, whereas GSS is far from it.

However, that said, for free software projects that wants to support Kerberos 5, we do acknowledge that no other framework provides a more portable and interoperable interface into the Kerberos 5 system. If your project needs to use Kerberos 5 specifically, we do recommend you to use GSS instead of the Kerberos 5 implementation specific APIs.

Appendix B Copying Information

B.1 GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ASCII without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text. The “publisher” means any person or entity that distributes copies of the Document to the public. A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING

You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3.

You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY

If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects.

If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages.

If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public.

It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS

You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version:

A. Use in the Title Page (and on the covers, if any) a title distinct from that of the Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Concept Index 71

Concept Index

A

AIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

C

command line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 Compiling your application . . . . . . . . . . . . . . . . . . . . . 8 Contributing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

D

Debian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2, 3 Download . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

F

FDL, GNU Free Documentation License . . . . . . . 63 FreeBSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Future goals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

H

Hacking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Header files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 HP-UX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

I

Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 invoking gss. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 IRIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

M

Mandrake . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 mechanism status codes . . . . . . . . . . . . . . . . . . . . . . . 15 Memory allocation failure . . . . . . . . . . . . . . . . . . . . . . . 8 Motorola Coldfire . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

N

NetBSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

O

OpenBSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Out of Memory handling. . . . . . . . . . . . . . . . . . . . . . . . 8

R

RedHat. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 RedHat Advanced Server . . . . . . . . . . . . . . . . . . . . . . . 3 Reporting Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

S

Solaris. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 status codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 SuSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 SuSE Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

T

Todo list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Tru64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

U

uClibc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 uClinux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

W

Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

API Index 72

API Index

gss. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 gss_accept_sec_context . . . . . . . . . . . . . . . . . . . . . . 31 gss_acquire_cred. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 gss_add_cred . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 gss_add_oid_set_member . . . . . . . . . . . . . . . . . . . . . . 50 GSS_CALLING_ERROR. . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 gss_canonicalize_name . . . . . . . . . . . . . . . . . . . . . . . 48 gss_check_version. . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 gss_compare_name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 gss_context_time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

gss_create_empty_oid_set . . . . . . . . . . . . . . . . . . . 53

gss_decapsulate_token . . . . . . . . . . . . . . . . . . . . . . . 54 gss_delete_sec_context . . . . . . . . . . . . . . . . . . . . . . 35 gss_display_name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 gss_display_status . . . . . . . . . . . . . . . . . . . . . . . . . . 51 gss_duplicate_name . . . . . . . . . . . . . . . . . . . . . . . . . . 49 gss_encapsulate_token . . . . . . . . . . . . . . . . . . . . . . . 54 GSS_ERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 gss_export_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 gss_export_sec_context . . . . . . . . . . . . . . . . . . . . . . 40 gss_get_mic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 gss_import_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 gss_import_sec_context . . . . . . . . . . . . . . . . . . . . . . 41 gss_indicate_mechs . . . . . . . . . . . . . . . . . . . . . . . . . . 52

gss_init_sec_context . . . . . . . . . . . . . . . . . . . . . . . . 25
gss_inquire_context . . . . . . . . . . . . . . . . . . . . . . . . . 37
gss_inquire_cred. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

                                                gss_inquire_cred_by_mech . . . . . . . . . . . . . . . . . . . 23
                                                gss_inquire_mech_for_saslname . . . . . . . . . . . . . . 55
                                                gss_inquire_mechs_for_name . . . . . . . . . . . . . . . . . 48
                                                gss_inquire_names_for_mech . . . . . . . . . . . . . . . . . 47
                                                gss_inquire_saslname_for_mech . . . . . . . . . . . . . . 55

gss_oid_equal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

                                                gss_process_context_token . . . . . . . . . . . . . . . . . . 36

gss_release_buffer . . . . . . . . . . . . . . . . . . . . . . . . . . 52
gss_release_cred. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24
gss_release_name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47
gss_release_oid_set . . . . . . . . . . . . . . . . . . . . . . . . . 53
GSS_ROUTINE_ERROR. . . . . . . . . . . . . . . . . . . . . . . . . . . . 17
GSS_S_... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16
GSS_SUPPLEMENTARY_INFO . . . . . . . . . . . . . . . . . . . . . . 17

                                                gss_test_oid_set_member. . . . . . . . . . . . . . . . . . . . . 53

gss_unwrap. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44
gss_userok. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57
gss_verify_mic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42
gss_wrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43
gss_wrap_size_limit . . . . . . . . . . . . . . . . . . . . . . . . . 39